### v1

# Out of Style: RAG’s Fragility to Linguistic Variation

Abstract
Despite the impressive performance of Retrieval-augmented Generation (RAG) systems across various NLP benchmarks, their robustness in handling real-world user-LLM interaction queries remains largely underexplored. This presents a critical gap for practical deployment, where user queries exhibit greater linguistic variations and can trigger cascading errors across interdependent RAG components. In this work, we systematically analyze how varying four linguistic dimensions (formality, readability, politeness, and grammatical correctness) impact RAG performance. We evaluate two retrieval models and nine LLMs, ranging from 3 to 72 billion parameters, across four information-seeking Question Answering (QA) datasets. Our results reveal that linguistic reformulations significantly impact both retrieval and generation stages, leading to a relative performance drop of up to 40.41% in Recall@5 scores for less formal queries and 38.86% in answer match scores for queries containing grammatical errors. Notably, RAG systems exhibit greater sensitivity to such variations compared to LLM-only generations, highlighting their vulnerability to error propagation due to linguistic shifts. These findings highlight the need for improved robustness techniques to enhance reliability in diverse user interactions.  [^1] Code is available at https://github.com/Springcty/RAG-fragility-to-linguistic-variation.
1 Introduction
Retrieval-augmented Generation (RAG) systems enhance Large Language Models (LLMs) by integrating external knowledge retrieval, grounding their output in factual context to improve accuracy and reliability (Lewis et al., 2021; Gao et al., 2024). However, their widespread integration into real-world applications (K2view, 2024) introduce potential challenges regarding robustness to linguistic variations. Users bring varied backgrounds, domains, and cultural contexts that naturally produce linguistic differences in their queries (Park et al., 2024; Li et al., 2020; Lorenzo-Dus & Bou-Franch, 2013). As Figure 1 illustrates, different from the carefully curated queries from traditional NLP benchmarks, real-world user-LLM queries tend to be less formal and frequently contain grammatical inconsistencies (Ouyang et al., 2023). Failing to account for these linguistic variations risks excluding a broad segment of users from effective interaction, especially for users whose linguistic expressions fall outside the narrow patterns these systems are tuned on (Liang et al., 2023). Moreover, unlike standalone LLMs, RAG systems incorporate multiple interdependent components, making them susceptible to cascading errors arising at both the retrieval and generation stages (Asai et al., 2023; Yoran et al., 2024a; Kim et al., 2025). A truly robust RAG system should maintain consistent retrieval effectiveness and generation quality across the full spectrum of user linguistic variations.
Figure 1: RAG systems demonstrate overall performance degradation when queries are rewritten to be less formal, more polite, less readable, and with grammatical errors. For the traditional NLP query (left), the RAG systems successfully retrieve related information and generate the correct answer, while the less formal queries (right) retrieve incorrect information. The linguistic variation on formality causes significant performance drops: 40.41% decrease in Recall@5 and 33.38% decrease in answer match (AM) score on the MS MARCO dataset.
We present a large-scale systematic investigation of how variations in linguistic characteristics affect the robustness of RAG systems. We focus on formality, readability, politeness, and grammatical correctness - fundamental aspects of linguistic variations that are critical and common in real-world user-LLM queries (Li et al., 2024; Yin et al., 2024; Hagen et al., 2024). By automatically rewriting queries across these dimensions, we analyze how linguistic variations impact each RAG system component, as well as potential cascading errors throughout the pipeline. Our evaluation encompasses two retrieval models, namely Contriever (Izacard et al., 2021) and ModernBERT (Nussbaum et al., 2024), and nine LLMs from three families (Llama 3.1, Qwen 2.5, Gemma 2) of varying scales across four open-domain Question Answering (QA) datasets: PopQA (Mallen et al., 2023), EntityQuestions (Sciavolino et al., 2022), MS MARCO (Bajaj et al., 2018), and Natural Questions (Kwiatkowski et al., 2019a).
Our experiments reveal significant vulnerabilities in RAG systems to linguistic variations. Retrieval analysis shows an average 15% relative performance degradation across all datasets and linguistic dimensions, with grammatical modifications most severely impacting recall while politeness variations show minimal effect. Generation analysis demonstrates average decreases of 16.52% (AM), 41.15% (EM), and 19.60% (F1) across all experimental settings. Notably, increasing the LLM scales doesn’t always help mitigate these performance gaps.
Furthermore, RAG systems exhibit greater vulnerability to linguistic variations than LLM-only generations, suggesting cascading errors across components. The PopQA dataset shows a 22.53% average performance drop in RAG systems versus only 10.78% in LLM-only generations. These findings highlight the urgent need to improve the robustness of retrieval components when handling linguistically varied queries. We also find that while advanced RAG techniques such as query expansion with HyDE (Gao et al., 2022) and documents reranking tend to improve overall performance, they remain similarly vulnerable to linguistic variations.
In summary, we present the first systematic analysis of the robustness of RAG systems to linguistic query reformulation. Our results demonstrate that despite strong performance on standard benchmarks, they remain fragile to inevitable real-world linguistic variations. These findings highlight the need for enhanced robustness techniques to improve reliability across diverse user interactions and inform design principles for next-generation RAG systems.
2 Related Work
Robustness of retrieval systems to linguistic variations.
Prior research investigated retrieval system robustness to noisy queries (Campos et al., 2023; Chen et al., 2022; 2023b) and specific linguistic variations including word substitutions (Wu et al., 2022), aspect changes and paraphrasing (Penha et al., 2022), typographical errors (Zhuang & Zuccon, 2021), and grammatical variations (Long et al., 2024). Our work provides the first holistic evaluation of RAG systems’ robustness to diverse linguistic variations and uncovers cascading failures across the entire RAG pipeline.
Robustness of language models to linguistic variations.
Prior research has examined impacts of syntactic perturbations Moradi & Samwald (2021); Singh et al. (2024), round-trip translation Bhandari & Chen (2023), politeness variations Yin et al. (2024), equivalent queries Cao et al. (2024) and scale of model size on robustness to grammatical errors (Yin et al., 2020; Hagen et al., 2024). Rawte et al. (2023) examined how formality and readability affect LLM performance in isolation. In contrast, our research investigates a broader spectrum of linguistic variations, generates data for each variation using LLM-based rewrites, and studies their compounding effects throughout the end-to-end RAG pipeline comprehensively.
Robustness of RAG systems.
While RAG systems demonstrate impressive performance (Lewis et al., 2021; Gao et al., 2024) and reduced hallucinations (Mallen et al., 2023), vulnerabilities exist with increasing retrieval context noise (Chen et al., 2023a), irrelevant contexts (Yoran et al., 2024b), and noise impacts (Fang et al., 2024). Yang et al. (2025) analyzed how spurious features affect RAG perform. These studies focus primarily on retrieved content noise rather than initial query variations. Our work uniquely demonstrates how diverse linguistic variations in user queries compound throughout the RAG pipeline, exposing critical vulnerabilities in systems serving diverse users.
3 Robustness Evaluation Approach
In our work, we explore the impact of the following linguistic aspects: Formality, Readability, Politeness and Grammatical Correctness - Round-Trip Translation and Typos. We explore these linguistic queries as they are essential dimensions of language variation that are prevalent and significant in real-world RAG interactions. We extend Rawte et al. (2023)’s findings on linguistic variations and LLM hallucinations by synthetically generating queries across four linguistic dimensions and four datasets to analyze each RAG pipeline component comprehensively. We first formulate our task, followed by defining each of our linguistic characteristics (Section 3.1), and elaborating on our query rewriting design (Section 3.2).
Task formulation.
Given a seed dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ and $y_{i}$ indicate $i$-th input and output, respectively, we reformulate each query $x_{i}\rightarrow x_{i}^{\prime}$ based on four linguistic aspects. A robust RAG system, composed of a retriever $\mathcal{R}$ and a generator $\mathcal{G}$ operating on corpus $\mathcal{C}$, should maintain performance when processing linguistically varied inputs. For the retrieval component, we expect retrieved documents $\mathbf{D}_{i}=\mathcal{R}(x_{i},\mathcal{C})$ and $\mathbf{D}_{i}^{\prime}=\mathcal{R}(x_{i}^{\prime},\mathcal{C})$ to both contain the answer $y_{i}$. For generation, given retrieved documents $\mathbf{D}_{i},\mathbf{D}_{i}^{\prime}$, a robust system should produce: $\mathcal{G}(x_{i},\mathbf{D}_{i})\approx\mathcal{G}(x_{i}^{\prime},\mathbf{D}_{i}^{\prime})\approx y_{i}$.
3.1 Linguistic Variations
Formality.
Formality in language lacks universal definition (Pavlick & Tetreault, 2016; Mosquera & Moreda, 2011; Fang & Cao, 2009), but encompasses situational factors (Hovy, 1987; Lahiri et al., 2011), grammar quality (Peterson et al., 2011), and specific linguistic elements like contractions (Heylighen et al., 1999). We quantify formality using the RoBERTa-based formality classifier from (Babakov et al., 2023b).
Readability.
Readability quantifies text comprehensibility through complexity and linguistic features. We employ the widely-used Flesch Reading Ease Score (FRES; Flesch 1948) to assess readability (Rawte et al., 2023; Han et al., 2024), defined in A.1.
Politeness.
Politeness is a sociocultural phenomenon defined as showing consideration of others (Wang, 2014). We calculate politeness scores using Polite Guard (Intel, 2024), an open-source NLP model from Intel that’s fine-tuned from BERT to classify text into four levels: polite, somewhat polite, neutral, and impolite.
Grammatical correctness.
In our work, we define grammatical correctness as the preservation of both grammaticality (Chomsky, 2002) and semantic fidelity. We alter the grammatical correctness through two approaches, inspired by Yin et al. (2020); Zhuang & Zuccon (2021); Lichtarge et al. (2019): (1) Typos, where random addition, deletion, or substitution operations are applied at a 20% probability per word, requiring an edit distance of at least 1; and (2) Round-trip translation (RTT) via English-Afrikaans-English using EasyNMT with opus-mt model, requiring the output to not be the exact same as the original query.
3.2 Query Rewriting
We systematically reformulate queries $x_{i}\rightarrow x_{i}^{\prime}$ across targeted linguistic dimensions while ensuring all rewrites satisfy dimension-specific thresholds and preserve semantic meaning (as specified in Appendix A.1). For each distinct dimension—formality, readability, and politeness—we take $5,000$ original queries and use GPT-4o-mini (OpenAI et al., 2024) to generate rewrites using three different prompts (detailed in Appendix G). We design distinct query sets across all datasets to be less formal, less readable, or more polite—directions chosen because the well-curated original datasets contained predominantly formal, readable, and neutral language, allowing our modifications to explore more realistic linguistic variations. This creates $15,000$ rewritten samples per dimension for each dataset, over which we average results to account for prompt stochasticity. For grammatical correctness, we similarly use $5,000$ original queries but apply deterministic transformations through typo introduction and round-trip translation as described in Section 3.1, creating $5,000$ modified samples for each approach. We ensure that grammatically modified queries differ from originals using GLEU and edit distance measurements. Example rewrites are shown in Table 1.
For each individual linguistic dimension and dataset, we construct controlled comparative datasets $\mathcal{D}^{\prime}=\{(x_{1}^{\prime},y_{1}),(x_{2}^{\prime},y_{2}),\ldots\}$. To validate our approach, we annotate $100$ queries across all linguistic variations, finding that 97.33% of the rewritten queries preserve the semantic meaning of the original query. Details are available in Appendix A.2.
Linguistic Dimension Example Rewrites
Politeness $\uparrow$ Original: complex carbohydrates are stored in animals in the form of
(MS MARCO) Rewritten: Would you be so kind as to share how complex carbohydrates are stored in animals?
Readability $\downarrow$ Original: who stars in the new movie the post
(Natural Questions) Rewritten: In the upcoming cinematic production titled “The Post,” which individuals have been cast in leading roles?
Formality $\downarrow$ Original: Who is the author of Dolores Claiborne?
(PopQA) Rewritten: Hey, do you happen to know who wrote Dolores Claiborne? I’m kinda curious!
Grammar $\downarrow$: Round-Trip Translation Original: Which company is HMS Blankney produced by?
(EntityQuestions) Rewritten: What company is producing HMS Blankey?
Grammar $\downarrow$: Typos Original: when did the japanese river otter become extinct
(Natural Questions) Rewritten: when did the japanese river otter ecome extinct
Table 1: Examples of query rewrites across different linguistic dimensions and datasets. Each pair shows the original query and its rewritten form.
4 Experimental Setups
4.1 Benchmarks
In this work, we use four open-domain QA datasets as seed datasets $\mathcal{D}$ and evaluate the effects how linguistic variations of the original queries affect RAG systems. PopQA (Mallen et al., 2023) is a large-scale entity-centric open-domain QA dataset about entities with a wide variety of popularity. EntityQuestions (Sciavolino et al., 2022) is a set of simple, entity-rich questions based on facts of Wikidata. MS MARCO (Bajaj et al., 2018) contains questions derived from real user search queries from Bing’s search logs. Natural Questions (Kwiatkowski et al., 2019a) contains questions consisting of real, anonymized, aggregated queries to the Google search engine. Both PopQA and EntityQuestions consist of well-structured, standardized, and simple queries, while queries from MS MARCO and Natural Questions exhibit free-form and arbitrary. We evaluate using PopQA’s test split, EntityQuestions’ dev split, and both dev and test splits from MS MARCO and Natural Questions.
4.2 Models
Retrieval.
We use two neural retrieval systems, namely Contriever (facebook/contriever;  Izacard et al. 2022) and ModernBERT Embed (nomic-ai/modernbert-embed-baseWarner et al. 2024; Nussbaum et al. 2024). Contriever is an unsupervised dense retriever built from BERT base architecture and is pre-trained using a contrastive learning framework. ModernBERT Embed is an embedding model trained from ModernBERT-base (Warner et al., 2024), bringing the new advances of ModernBERT to embeddings. We use the retrieval system implementation by  Izacard et al. (2022) for both retrieval models. [^2] https://github.com/facebookresearch/contriever
Generation.
We evaluate nine most advanced open-source LLMs from three model families with various scales: Llama 3.1 (Llama-3.1-8B-Instruct, Llama-3.1-70B-InstructGrattafiori et al. 2024), Qwen 2.5 (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct;  Yang et al. 2024; Team 2024), and Gemma 2 (gemma-2-2b-it, gemma-2-9b-it, gemma-2-27b-itTeam et al. 2024). Detailed hyperparameter settings are included in the Appendix B.1.
We use few-shot prompting to ensure that the model outputs are in the correct format. For each dataset and linguistic characteristic, we include two question-answer pairs in the context: one random original query with its answer, and its corresponding linguistically rewritten version with the same answer. This balanced approach exposes the model to both original and rewritten query formats to ensure fairness. We also include the top five retrieved passages in the context. Detailed prompts are provided in Appendix H.
4.3 Metrics
Retrieval.
We employ Recall@k (R@k) (Karpukhin et al., 2020), which calculates the fraction of the retrieved documents containing gold answers. We use $k=5$ as our primary setup, and evaluate the effect of varying $k$ in our analysis.
Generation.
The generation stage is assessed using a comprehensive set of metrics: Answer Match (AM) measures the percentage of the predictions that any substring of the prediction is an exact match of any of the ground truth answers, Exact Match (EM) measures the percentage of predictions that exactly match any of the ground truth answers, and F1 Score (F1) captures the harmonic mean of precision and recall in generated responses.
5 RAG Robustness Experimental Results
In this section, we progress from component-level to end-to-end RAG system analysis, followed by an assessment of advanced techniques (query expansion and re-ranking) and their ability to mitigate performance drops when faced with linguistic variations.
5.1 Retrieval Analysis
We conduct a comprehensive analysis of retrieval systems performance across two candidate retrievers: Contriever and ModernBERT Embed. The results are shown in Table 2.
Query variations based on linguistic dimensions degrade retrieval performance.
Our analysis (Table 2) reveals significant linguistic fragility in retrieval systems, with performance degradation averaging 16.7% for Contriever and 13.3% for ModernBERT across all modifications. Results show highest sensitivity on PopQA (19.78% average impact), particularly to grammatical transformations (29.34% from RTT). MS MARCO exhibits the second-highest impact (16.78%), with striking sensitivity to formality changes (40.41% with Contriever), suggesting retrieval systems may be implicitly optimized for specific linguistic patterns, limiting effectiveness when handling diverse query variations.
Linguistics Retriever PopQA Entity MARCO NQ $\Delta$ Q-len
Readability Contriever 18.45 (0.61) 8.61 (0.64) 21.10 (0.25) 7.69 (0.60) 5.23
ModernBERT 17.73 (0.65) 13.17 (0.61) 14.58 (0.40) 10.08 (0.65)
Gram. (RTT) Contriever 29.00 (0.59) 14.57 (0.68) 9.14 (0.34) 23.50 (0.62) -0.26
ModernBERT 29.68 (0.62) 14.85 (0.66) 17.54 (0.32) 18.24 (0.67)
Gram. (Typos) Contriever 27.79 (0.59) 14.80 (0.68) 15.53 (0.34) 30.83 (0.62) 0.02
ModernBERT 22.48 (0.62) 11.01 (0.66) 12.30 (0.32) 13.45 (0.67)
Formality Contriever 19.96 (0.70) 10.71 (0.68) 40.41 (0.25) 15.35 (0.65) 13.65
ModernBERT 13.67 (0.74) 8.05 (0.68) 15.55 (0.40) 9.51 (0.69)
Politeness Contriever 8.30 (0.62) 1.70 (0.67) 16.44 (0.26) 1.16 (0.60) 7.29
ModernBERT 10.70 (0.67) 3.39 (0.67) 5.18 (0.40) 4.97 (0.65)
Table 2: Relative retrieval performance drop (%) in R@5 on rewritten queries across datasets. (Original scores) shown in gray parentheses. Bold indicates the largest degradation value per retriever. $\Delta$ Q-len represents average query token-length change. Query linguistic variations degrades retrieval performance consistently across all linguistic characteristics.
Grammatical variations have the highest impact on retrieval performance.
On average, grammaticality rewrites emerge as the most impactful linguistic variation on recall performance. Round-trip translation degrades recall by an average of 19.56% across all datasets. Interestingly, ModernBERT shows greater vulnerability to these structural transformations (20.12% drop) compared to Contriever (19% drop). Typographical errors present another significant challenge, causing an average recall reduction of 18.51%. However, the retrievers display opposite behavior patterns with typos: Contriever exhibits substantially lower robustness (22.22% drop) than ModernBERT (14.81% drop). This suggests that ModernBERT’s diverse training data mixture likely enables it to develop greater robustness to character-level grammatical perturbations compared to Contriever.
Politeness variations have minimal impact on retrieval performance.
Politeness variations have the least impact on retrieval performance, with an average recall drop of only 6.48% across all datasets and retrievers. This stands in stark contrast to grammatical variations (19.56%) and typos (18.51%). The minimal effect is most evident in Natural Questions with Contriever (1.16%) and EntityQuestions with Contriever (1.70%). This suggests that retrieval models effectively filter out social courtesy markers while preserving their focus on the query’s core semantic content and keywords, maintaining robust performance despite changes in query politeness level.
Retrieval performance drops independent of query length.
Our analysis, as shown in Table 2 demonstrates that query length changes do not directly correlate with retrieval performance. Queries with increased formality showed substantial length increases (+13.65 tokens) yet produced inconsistent performance impacts across datasets. Conversely, round-trip translated queries were marginally shorter (-0.26 tokens) but consistently caused significant performance degradation. This indicates retrieval models respond more to linguistic quality (grammatical correctness, readability) than to query length itself, highlighting the need for systems robust to linguistic variations rather than optimized for specific query lengths.
Scaling up number of documents improves performance.
Table 2 shows performance degradation ($\Delta$R@K) decreasing as $K$ increases, indicating linguistic perturbations cause relevant documents to slide down rather than disappear from the ranked list. As an example, for rewrites with typos, $\Delta$ R@K for Contriever decreases from 22.2 at R@5 to 12.0 at R@100, and for ModernBERT from 20.1 to 9.8. This is detailed further in Appendix C.1. This ranking deterioration forces downstream language models to operate with suboptimal information, potentially compromising response quality. We further investigate this hypothesis in Section 5.4.2 by examining if rerankers can improve recall scores for Top-5 retrieved documents.
5.2 Generation Analysis
The RAG experiment results on ModernBERT retriever and nine LLMs are presented in Table 3. Overall, RAG systems show performance degradation on all linguistic variations.
Model Rew. $\downarrow$ (Ori.) Readability Grammatical Correctness
PopQA NQ MARCO Entity PopQA NQ MARCO Entity
RTT Typos RTT Typos RTT Typos RTT Typos
gemma-2-2b-it 20.71 (0.52) 16.82 (0.45) 17.44 (0.20) 16.43 (0.51) 32.73 20.26 (0.49) 28.09 10.44 (0.43) 29.81 10.19 (0.16) 22.47 10.36 (0.54)
gemma-2-9b-it 17.12 (0.54) 13.95 (0.49) 11.13 (0.20) 15.25 (0.54) 33.90 19.60 (0.51) 24.50 8.19 (0.48) 25.69 6.98 (0.16) 22.92 10.39 (0.56)
gemma-2-27b-it 16.15 (0.56) 11.33 (0.52) 8.33 (0.20) 12.68 (0.55) 32.81 18.21 (0.53) 24.03 6.81 (0.51) 27.71 3.79 (0.15) 20.73 10.69 (0.58)
Llama-3.1-8B-Instruct 17.27 (0.55) 15.07 (0.52) 15.23 (0.22) 13.30 (0.52) 35.27 21.95 (0.50) 27.45 10.54 (0.50) 32.37 12.37 (0.19) 25.06 12.53 (0.56)
Llama-3.1-70B-Instruct 17.66 (0.58) 15.77 (0.54) 13.30 (0.22) 16.09 (0.55) 35.30 19.49 (0.54) 25.43 8.80 (0.52) 24.83 5.38 (0.17) 21.61 12.71 (0.57)
Qwen2.5-3B-Instruct 22.35 (0.47) 18.65 (0.42) 13.27 (0.21) 24.14 (0.46) 32.89 23.77 (0.44) 30.74 15.65 (0.44) 32.14 12.31 (0.17) 24.53 13.20 (0.49)
Qwen2.5-7B-Instruct 19.50 (0.53) 17.15 (0.48) 11.55 (0.21) 20.35 (0.51) 34.06 22.67 (0.49) 29.67 13.19 (0.49) 24.28 6.64 (0.17) 21.77 11.99 (0.55)
Qwen2.5-32B-Instruct 17.00 (0.56) 14.01 (0.52) 32.10 (0.28) 18.41 (0.53) 34.11 19.69 (0.52) 24.80 9.28 (0.52) 26.67 7.31 (0.16) 20.96 11.62 (0.57)
Qwen2.5-72B-Instruct 16.25 (0.56) 13.62 (0.54) 8.36 (0.20) 17.03 (0.54) 33.68 19.27 (0.53) 24.28 7.74 (0.53) 27.47 8.22 (0.16) 19.56 10.35 (0.58)
Avg 18.22 (0.54) 15.15 (0.50) 14.52 (0.22) 17.07 (0.52) 33.86 20.55 (0.51) 26.55 10.07 (0.49) 27.88 8.13 (0.17) 22.18 11.54 (0.56)
Model Rew. $\downarrow$ (Ori.) Formality Politeness PopQA NQ MARCO Entity PopQA NQ MARCO Entity gemma-2-2b-it 13.64 (0.61) 12.44 (0.46) 23.31 (0.17) 7.66 (0.60) 8.37 (0.52) 4.42 (0.44) 11.16 (0.20) 2.96 (0.59) gemma-2-9b-it 11.67 (0.64) 9.95 (0.50) 19.68 (0.17) 8.85 (0.62) 8.54 (0.54) 3.60 (0.48) 8.19 (0.20) 3.01 (0.61) gemma-2-27b-it 12.21 (0.65) 9.92 (0.52) 19.13 (0.17) 7.19 (0.63) 7.93 (0.56) 3.38 (0.51) 6.13 (0.20) 1.97 (0.61) Llama-3.1-8B-Instruct 12.23 (0.65) 10.51 (0.52) 17.59 (0.20) 7.16 (0.62) 7.59 (0.55) 3.46 (0.51) 9.90 (0.23) 3.50 (0.59) Llama-3.1-70B-Instruct 11.48 (0.67) 12.09 (0.54) 19.00 (0.20) 7.19 (0.64) 7.00 (0.57) 6.29 (0.53) 9.89 (0.23) 3.61 (0.61) Qwen2.5-3B-Instruct 10.97 (0.58) 11.38 (0.45) 20.87 (0.18) 7.51 (0.56) 11.54 (0.47) 5.91 (0.42) 12.30 (0.21) 7.74 (0.55) Qwen2.5-7B-Instruct 12.99 (0.63) 11.46 (0.50) 19.84 (0.18) 8.08 (0.61) 10.76 (0.54) 5.05 (0.48) 7.08 (0.21) 4.69 (0.60) Qwen2.5-32B-Instruct 11.34 (0.65) 7.85 (0.53) 14.80 (0.18) 6.24 (0.62) 8.65 (0.56) 4.76 (0.52) 6.66 (0.21) 5.64 (0.60) Qwen2.5-72B-Instruct 10.68 (0.66) 9.11 (0.55) 20.72 (0.18) 5.75 (0.63) 7.36 (0.56) 5.45 (0.54) 16.62 (0.20) 3.86 (0.62) Avg 11.91 (0.64) 10.52 (0.51) 19.44 (0.18) 7.29 (0.61) 8.64 (0.54) 4.70 (0.49) 9.77 (0.21) 4.11 (0.60)
Table 3: RAG performance on answer match (AM) scores using ModernBERT retriever with the Gemma 2, Llama 3.1, and Qwen 2.5 model families across four datasets. Results show relative percentage performance degradation on rewritten queries (Rew. % $\downarrow$) and the original query performance (Ori.) within parentheses in gray. For RTT, it has the same original scores as Typos. The largest degradation value among four datasets is in bold. All systems exhibit performance drops across all linguistic variations and datasets.
RAG systems are sensitive to linguistic variations.
As illustrated in Table 3, across all datasets, we observe a noticeable overall degradation in performance when queries are rewritten to become less formal, more polite, less readable, or have grammatical errors. Across all datasets, linguistic dimensions, and experimental settings, we found average drops of 16.52% (AM), 41.15% (EM), and 19.60% (F1). The PopQA dataset shows the highest sensitivity to all linguistic variations, with an average performance drop of 18.64% on AM scores. Particularly notable were the effects of reduced readability (18.22% degradation) and round-trip translation (33.86% degradation). These findings suggest that while the RAG systems perform well on standard NLP benchmarks with structured queries, they remain vulnerable to common linguistic variations. The full experiment results are shown in Appendix D.
Politeness reformulations yield different impacts on AM and EM scores.
When queries are rephrased to be more polite, we find that the AM scores remain relatively similar to those of the original, with less than 10% change for all datasets. However, there are significant drops in exact match (EM) scores. Specifically, we observe 44.57% and 18.32% drops in EM scores for queries from the Natural Questions and PopQA datasets, respectively. Through manual checks, we find that LLMs tend to generate more complete or formal responses in polite query formulations, which results in lower EM rates but similar performance on AM.
The round-trip translation errors and typos highlight different sensitivities.
Round-trip translation, which introduces structural sentence transformations, generally causes notable decreases across all datasets, showing 33.86%, 26.55%, 27.88%, and 22.18% drops in AM scores in the PopQA, Natural Questions, MS MARCO, and EntityQuestions, respectively. In contrast, typos, mainly introducing surface-level grammatical errors, produce moderate but less drastic performance degradation. This finding is consistent with the retrieval experiment results, suggesting that RAG systems are more vulnerable to structural transformations than superficial grammatical mistakes.
5.3 Retrieval Method and LLM Scale Influence
In this section, we are going to investigate the influence of different retrieval methods and LLM scales on the robustness of the RAG systems. The main results are shown in Figure 2.
Figure 2: PopQA and Natural Questions (NQ) LLMs scaling results, augmented with ModernBERT, Contriever, and LLM-only generation (Vanilla). Retrieval-augmented generation is more sensitive to linguistic variations than the LLM-only generation.
RAG systems with ModernBERT retrieval show greater robustness to linguistic variations.
As shown in Figure 2, generation results based on ModernBERT retrieval consistently outperform those with Contriever retrieval across both original and rewritten queries. Notably, RAG systems with ModernBERT demonstrate superior robustness to linguistic variations, exhibiting an average performance drop of only 19.52% on rewritten queries compared to 24.38% for Contriever. This suggests ModernBERT retrieval maintains better semantic understanding when handling linguistically varied inputs.
RAG systems show higher sensitivity to linguistic variations than LLM-only generations.
For PopQA, we observe an average performance drop of 22.53% across all linguistic variations and both retrieval models, while LLM-only generations experience only a 10.78% reduction. Even more striking, Figure 2 shows that there is barely any performance difference in LLM-only generation on formality rewrites, which suggests that errors are cascaded from the retrieval component to the generation component in the RAG system. These findings further indicate that retrieval components represent the primary vulnerability in RAG systems when handling linguistic variations.
LLM scaling doesn’t always help with mitigating performance gaps in RAG systems.
Notably, the performance gap between original and rewritten queries narrows for formality and readability variations as LLMs scale up (see Figure 2). Specifically, PopQA shows reduced degradation on less readable queries from 22.35% at 3B to 16.25% at 72B parameter. This suggests that larger models can extract relevant information better from retrieved contexts. However, this scaling benefit remains selective and limited; for round-trip translation variations, the performance gap actually widens with increased model size. This counterintuitive finding may be attributed to the structural transformations introduced during translation that become more problematic for larger models attempting more precise reasoning.
5.4 Exploring the Robustness of Advanced RAG Systems
Figure 3: Retrieval (ModernBERT, R@5 Score) and generation (Qwen2.5-7B-Instruct, AM Score) performance across different RAG settings on PopQA. We find that (1) adding HyDE and Rerank to the RAG pipeline improves the robustness to linguistic variations, but still lags behind original queries in performance. (2) HyDE improves robustness but slightly reduces performance on original queries. (3) Reranking improves performance on both original and rewritten queries.
Many modern RAG systems include more components than simply retrieval and generation which aim to make them more useful for users (Gao et al., 2022). In our experiments, we broadly explore whether the addition of two advanced components—Query Expansion and Reranking—improves RAG robustness. Detailed results of the experiment can be found in Appendices E and F.
5.4.1 Query Expansion
We evaluate Hypothetical Document Embeddings (HyDE; Gao et al. 2022) for query expansion. Figure 3 reveals that HyDE improves ModernBERT’s retrieval on linguistically varied queries (readability, typos, formality, politeness) by 2.61% on average, but severely impairs performance on round-trip translated queries (11% decrease). For original queries, HyDE consistently reduces ModernBERT’s retrieval effectiveness by 5.43%. Similarly, generation quality increases for rewritten queries across PopQA (3.77%) but decreases for original queries (1.92%). These findings suggest HyDE provides insufficient benefits for ModernBERT, underscoring the need for more effective query expansion methods.
5.4.2 Reranker
The retriever must be efficient for large document collections containing millions of entries, although it may sometimes retrieve irrelevant candidates. To address this, we incorporate a Cross-Encoder-based re-ranker to significantly enhance the quality of final answers. Specifically, we employ the MS MARCO Cross-Encoders developed by Reimers & Gurevych (2019) to re-rank passages retrieved by ModernBERT and Contriever. As illustrated in Figure 3, re-ranking substantially improves retrieval performance, particularly for rewritten queries, achieving an average improvement of 16.56% compared to only 7.40% for original queries. In contrast, generation results show a modest improvement of 1.83% for original queries and a more substantial improvement of 8.50% for rewritten queries. These findings suggest that the effectiveness of re-ranking is especially pronounced when handling rewritten queries and highlight the importance of improving the robustness of retrieval systems.
6 Conclusion
We conduct the first large-scale, systematic investigation into how linguistic variations—specifically formality, readability, politeness, and grammatical correctness—impact the robustness of RAG systems. Our analysis reveals that both the retrieval and generation components suffer performance degradation when faced with linguistic variations. Notably, RAG systems exhibit greater vulnerability to linguistic variations compared to LLM-only generations, indicating potential cascading errors within the retrieval-generation pipeline. Crucially, increasing the scale of LLMs does not consistently mitigate these robustness issues, and even advanced retrieval techniques like HyDE and reranking show similar susceptibility. These findings highlight the need to develop strategies that ensure reliable performance across linguistically varied queries, guiding future improvements of real-world RAG systems.
Appendix A Linguistic Rewriting Settings
A.1 Linguistic Variations
In this section, we detail our design of linguistic variations as well as our qualitative analysis of the queries we generated.
Query rewrite design
: For each linguistic dimension, we establish quantitative thresholds to ensure meaningful rewrites from original query $x_{i}$ to the rewritten query $x_{i}^{\prime}$:
  • Formality: $x_{i}^{\prime}$ must score below 0.5 probability using our formality classifier
  • Readability: $x_{i}^{\prime}$ must score below 60 on the Flesch Reading Ease Score
  • Politeness: The sum of somewhat polite and polite score on $x_{i}^{\prime}$ must above 0.5 using Polite Guard classification model.
  • Grammatical correctness (Typos): $x_{i}^{\prime}$ must have edit distance $\geq 1$ from $x_{i}$ and must not have the same GLEU score. Edit distance between two strings is the minimum number of operations required to transform one string into the other.
  • Grammatical correctness (Round-trip translation): $x_{i}^{\prime}$ must differ from $x_{i}$ after English-Afrikaans-English translation, constrained by ensuring that both queries do not have the same GLEU score and an edit distance of 0.
Each rewrite for each prompt must adhere to these thresholds in order to develop the final dataset we use. In addition to this, we want to ensure that the rewritten queries are semantically similar to the original query. Therefore, in the query-rewriting process, we add an additional constraint where the sentence similarity between the original and rewritten queries must be greater than 0.7, as assessed by the MPNet-v2 model Song et al. (2020) in the SentenceBert Library Reimers & Gurevych (2019).
Evaluation
: Flesch Reading score is defined as follows:
FRES=206.8351.015(total wordstotal sentences)84.6(total syllablestotal $\text{FRES}=206.835-1.015\left(\frac{\text{total words}}{\text{total sentences}}\right)-84.6\left(\frac{\text{total syllables}}{\text{total words}}\right)$ (1)
A.2 Semantic Preservation of Linguistic Variations
To ensure our linguistic rewrites preserve semantic meaning, we conducted a qualitative evaluation where two researchers annotated 20 randomly sampled query rewrites per linguistic property for the EntityQuestions dataset, totaling 100 queries examined by both annotators. Our primary assessment criterion was whether the rewritten question remained semantically similar enough to warrant the same answer as the original query.
For typos and back-translation variations specifically, we established a threshold of maximum 2 character edits on entities in the questions to determine if a query remained semantically equivalent to its original form. This threshold helps maintain a balance between introducing realistic linguistic variations and preserving essential meaning.
Our analysis revealed high semantic preservation rates across all linguistic dimensions:
Linguistic Dimension Semantic Preservation (%) Inter-annotator Agreement (%)
Politeness 100.0 100.0
Formality 98.6 100.0
Readability 96.7 96.7
Typos 95.0 100.0
Round-trip Translation 95.0 100.0
Table 4: Percentage of queries that maintain semantic similarity after linguistic rewrites and inter-annotator agreement rates.
These results demonstrate the robustness of our rewriting methodology in preserving semantic content while introducing natural linguistic variations. The high inter-annotator agreement further validates the reliability of our assessment approach. This semantic preservation is critical for our experimental design, as it ensures that performance differences observed in RAG systems can be attributed to linguistic variations rather than semantic drift.
Appendix B Detailed RAG Experimental Settings
B.1 LLMs Generation Hyperparameter Setting
We set temperature as 0.5 and top_q 0.90 to strike a balance between creativity and accuracy. The max_tokens is set to 128 considering the gold answers length.
Appendix C Full Retrieval Experiment Results
In this section we provide the absolute results from Contriever retriever experiments that led to Table 2
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
EntityQuestions RTT 0.6838 0.7332 0.7710 0.8422
EntityQuestions Typos 0.6838 0.7332 0.7710 0.8422
EntityQuestions Formality 0.6846 0.7292 0.7616 0.8240
EntityQuestions Politeness 0.6744 0.7218 0.7594 0.8310
EntityQuestions Readability 0.6434 0.6994 0.7452 0.8354
MS MARCO RTT 0.3412 0.4068 0.4694 0.6102
MS MARCO Typos 0.3412 0.4068 0.4694 0.6102
MS MARCO Formality 0.2534 0.3252 0.4048 0.5598
MS MARCO Politeness 0.2620 0.3414 0.4152 0.5664
MS MARCO Readability 0.2512 0.3280 0.4072 0.5718
Natural Questions RTT 0.6246 0.7118 0.7834 0.8822
Natural Questions Typos 0.6246 0.7118 0.7834 0.8822
Natural Questions Formality 0.6472 0.7372 0.7954 0.8868
Natural Questions Politeness 0.6012 0.6946 0.7588 0.8488
Natural Questions Readability 0.5978 0.6882 0.7582 0.8536
PopQA RTT 0.5938 0.6614 0.7148 0.8192
PopQA Typos 0.5938 0.6614 0.7148 0.8192
PopQA Formality 0.6974 0.7574 0.8066 0.8760
PopQA Politeness 0.6220 0.6942 0.7576 0.8534
PopQA Readability 0.6108 0.6856 0.7368 0.8344
Table 5: Contriever Retrieval performance (R@k) of original queries across datasets and linguistic modifications.
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
EntityQuestions RTT 0.5842 0.6428 0.6926 0.7896
EntityQuestions Typos 0.5826 0.6452 0.6964 0.7836
EntityQuestions Formality 0.6113 0.6626 0.7025 0.7815
EntityQuestions Politeness 0.6629 0.7090 0.7507 0.8253
EntityQuestions Readability 0.5887 0.6471 0.6979 0.7935
MS MARCO RTT 0.3100 0.3662 0.4328 0.5754
MS MARCO Typos 0.2882 0.3460 0.4052 0.5440
MS MARCO Formality 0.1502 0.2020 0.2722 0.4365
MS MARCO Politeness 0.2189 0.2933 0.3764 0.5448
MS MARCO Readability 0.1982 0.2716 0.3467 0.5235
Natural Questions RTT 0.4778 0.5718 0.6506 0.7898
Natural Questions Typos 0.4320 0.5230 0.6180 0.7730
Natural Questions Formality 0.5479 0.6440 0.7215 0.8456
Natural Questions Politeness 0.5942 0.6832 0.7516 0.8443
Natural Questions Readability 0.5519 0.6481 0.7253 0.8346
PopQA RTT 0.4216 0.4854 0.5422 0.6654
PopQA Typos 0.4288 0.4890 0.5482 0.6724
PopQA Formality 0.5582 0.6269 0.6800 0.7856
PopQA Politeness 0.5704 0.6426 0.7045 0.8081
PopQA Readability 0.4981 0.5635 0.6192 0.7332
Table 6: Contriever Retrieval performance (R@k) of rewritten queries across datasets and linguistic modifications
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
EntityQuestions RTT 0.6614 0.7184 0.7598 0.8284
EntityQuestions Typos 0.6614 0.7184 0.7598 0.8284
EntityQuestions Formality 0.6798 0.7240 0.7558 0.8150
EntityQuestions Politeness 0.6730 0.7214 0.7594 0.8276
EntityQuestions Readability 0.6132 0.6758 0.7254 0.8108
MS MARCO RTT 0.3204 0.3916 0.4574 0.5680
MS MARCO Typos 0.3204 0.3916 0.4574 0.5680
MS MARCO Readability 0.3982 0.4818 0.5604 0.6746
MS MARCO Formality 0.4074 0.4896 0.5644 0.6720
MS MARCO Politeness 0.4030 0.4840 0.5552 0.6638
Natural Questions RTT 0.6690 0.7556 0.8110 0.8878
Natural Questions Typos 0.6690 0.7556 0.8110 0.8878
Natural Questions Readability 0.6512 0.7300 0.7872 0.8614
Natural Questions Formality 0.6874 0.7700 0.8246 0.8944
Natural Questions Politeness 0.6538 0.7326 0.7860 0.8600
PopQA RTT 0.6280 0.6952 0.7508 0.8344
PopQA Typos 0.6280 0.6952 0.7508 0.8344
PopQA Readability 0.6518 0.7168 0.7682 0.8432
PopQA Formality 0.7408 0.7922 0.8316 0.8832
PopQA Politeness 0.6744 0.7418 0.7962 0.8688
Table 7: ModernBERT Retrieval performance (R@k) for original queries across datasets and linguistic modifications
Dataset Linguistics R@5 R@10 R@20 R@100
PopQA Readability 0.5363 0.6148 0.6751 0.7796
PopQA RTT 0.4416 0.5090 0.5668 0.6856
PopQA Typos 0.4868 0.5646 0.6242 0.7406
PopQA Formality 0.6395 0.7109 0.7649 0.8493
PopQA Politeness 0.6023 0.6776 0.7334 0.8317
EntityQuestions Readability 0.5325 0.5992 0.6607 0.7729
EntityQuestions RTT 0.5632 0.6260 0.6784 0.7800
EntityQuestions Typos 0.5886 0.6518 0.7038 0.7918
EntityQuestions Formality 0.6251 0.6739 0.7156 0.7949
EntityQuestions Politeness 0.6502 0.7002 0.7436 0.8199
MS MARCO Readability 0.3401 0.4291 0.5119 0.6461
MS MARCO RTT 0.2642 0.3340 0.3908 0.5146
MS MARCO Typos 0.2810 0.3540 0.4188 0.5396
MS MARCO Formality 0.3441 0.4311 0.5113 0.6479
MS MARCO Politeness 0.3821 0.4637 0.5385 0.6579
Natural Questions Readability 0.5855 0.6741 0.7435 0.8393
Natural Questions RTT 0.5470 0.6416 0.7192 0.8328
Natural Questions Typos 0.5790 0.6838 0.7512 0.8502
Natural Questions Formality 0.6220 0.7177 0.7890 0.8762
Natural Questions Politeness 0.6213 0.7054 0.7701 0.8539
Table 8: ModernBERT Retrieval performance (R@k) for rewritten queries across datasets and linguistic modifications
C.1 Scaling Number of documents
Figure 4: Average Recall@K increase as Number of Top-K Documents increases – Contriever.
Figure 5: Average Recall@K increase as Number of Top-K Documents increases – ModernBERT.
As mentioned in Section 5.1, the scaling of the number of documents decreases the degradation in performance, but does not mitigate the overall issue. As you can see in Figures 4 and 5, as K increases, higher recall benefits both original and rewritten queries, with performance gaps narrowing as correctly ranked documents appear at lower positions—suggesting linguistic variations primarily affect ranking order rather than complete retrieval failure. This hypothesis is confirmed in Section 5.4.2, where the reranking shows considerable performance gains at R@5, showing that retrieval systems tend to push the correct documents for rewritten queries to lower ranks, and reranking helps prioritize them again, which is demonstrated by the reduction in performance degradation in Figure 3.
Appendix D Full RAG Experiment Results
Contriever, Formality, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.5858 0.4719 19.45% 0.4432 0.3742 15.57% 0.1416 0.0955 32.58% 0.6026 0.5446 9.62%
gemma-2-9b-it 0.6006 0.4995 16.83% 0.4784 0.4177 12.68% 0.1420 0.0971 31.64% 0.6148 0.5509 10.39%
gemma-2-27b-it 0.6126 0.5094 16.85% 0.5046 0.4447 11.88% 0.1384 0.1025 25.96% 0.6300 0.5691 9.67%
Llama-3.1-8B-Instruct 0.6122 0.5098 16.73% 0.4930 0.4376 11.24% 0.1580 0.1121 29.03% 0.6156 0.5621 8.70%
Llama-3.1-70B-Instruct 0.6386 0.5297 17.05% 0.5170 0.4331 16.22% 0.1586 0.1057 33.38% 0.6352 0.5747 9.52%
Qwen2.5-3B-Instruct 0.5648 0.4596 18.63% 0.4288 0.3669 14.43% 0.1416 0.1002 29.24% 0.5626 0.5086 9.60%
Qwen2.5-7B-Instruct 0.5910 0.4864 17.70% 0.4774 0.4041 15.35% 0.1544 0.1085 29.71% 0.6106 0.5529 9.44%
Qwen2.5-32B-Instruct 0.6176 0.5128 16.97% 0.5164 0.4586 11.19% 0.1494 0.1115 25.39% 0.6242 0.5720 8.36%
Qwen2.5-72B-Instruct 0.6266 0.5249 16.24% 0.5280 0.4709 10.82% 0.1462 0.1079 26.22% 0.6366 0.5872 7.76%
Contriever, Politeness, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5206 0.4903 5.83% 0.4870 0.4857 0.27% 0.1696 0.1463 13.72% 0.6300 0.5691 9.67%
gemma-2-2b-it 0.4798 0.4508 6.04% 0.4040 0.4147 2.66% 0.1724 0.1400 18.79% 0.6026 0.5446 9.62%
gemma-2-9b-it 0.4974 0.4685 5.80% 0.4560 0.4528 0.70% 0.1736 0.1487 14.36% 0.6148 0.5509 10.39%
Llama-3.1-70B-Instruct 0.5384 0.5059 6.03% 0.4966 0.4811 3.13% 0.1900 0.1528 19.58% 0.6352 0.5747 9.52%
Llama-3.1-8B-Instruct 0.5142 0.4809 6.47% 0.4806 0.4735 1.48% 0.1874 0.1593 15.01% 0.6156 0.5621 8.70%
Qwen2.5-32B-Instruct 0.5184 0.4799 7.42% 0.4948 0.4787 3.25% 0.1776 0.1525 14.15% 0.6242 0.5720 8.36%
Qwen2.5-3B-Instruct 0.4436 0.4026 9.24% 0.3938 0.3830 2.74% 0.1680 0.1423 15.28% 0.5626 0.5086 9.60%
Qwen2.5-72B-Instruct 0.5216 0.4906 5.94% 0.5156 0.4991 3.21% 0.1722 0.1348 21.72% 0.6366 0.5872 7.76%
Qwen2.5-7B-Instruct 0.4830 0.4544 5.92% 0.4402 0.4329 1.67% 0.1814 0.1552 14.44% 0.6106 0.5529 9.44%
Contriever, Readability, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5252 0.4259 18.91% 0.4936 0.4499 8.85% 0.1668 0.1512 9.35% 0.5600 0.5064 9.57%
gemma-2-2b-it 0.4778 0.3762 21.26% 0.4146 0.3559 14.15% 0.1686 0.1413 16.21% 0.5254 0.4660 11.31%
gemma-2-9b-it 0.5026 0.4095 18.53% 0.4602 0.4041 12.20% 0.1638 0.1439 12.13% 0.5488 0.4872 11.22%
Llama-3.1-70B-Instruct 0.5426 0.4414 18.65% 0.5068 0.4383 13.51% 0.1780 0.1482 16.74% 0.5636 0.4947 12.23%
Llama-3.1-8B-Instruct 0.5184 0.4186 19.25% 0.4880 0.4235 13.22% 0.1798 0.1512 15.91% 0.5388 0.4851 9.97%
Qwen2.5-32B-Instruct 0.5184 0.4168 19.60% 0.4964 0.4312 13.13% 0.1794 0.1607 10.41% 0.5442 0.4725 13.18%
Qwen2.5-3B-Instruct 0.4480 0.3379 24.57% 0.3880 0.3272 15.67% 0.1654 0.1371 17.13% 0.4762 0.3840 19.36%
Qwen2.5-72B-Instruct 0.5272 0.4310 18.25% 0.5140 0.4523 12.01% 0.1714 0.1551 9.49% 0.5548 0.4816 13.19%
Qwen2.5-7B-Instruct 0.4864 0.3890 20.02% 0.4418 0.3817 13.60% 0.1770 0.1523 13.97% 0.5154 0.4412 14.40%
Contriever, Round-trip Translation, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.4942 0.3290 33.43% 0.4808 0.3438 28.49% 0.1270 0.0952 25.04% 0.5914 0.4744 19.78%
gemma-2-2b-it 0.4628 0.3088 33.28% 0.4090 0.2700 33.99% 0.1282 0.0856 33.23% 0.5568 0.4378 21.37%
gemma-2-9b-it 0.4746 0.3148 33.67% 0.4512 0.3208 28.90% 0.1304 0.0970 25.61% 0.5672 0.4432 21.86%
Llama-3.1-70B-Instruct 0.5064 0.3288 35.07% 0.1420 0.3276 130.70% 0.1420 0.1050 26.06% 0.5818 0.4606 20.83%
Llama-3.1-8B-Instruct 0.4708 0.3100 34.15% 0.4622 0.3184 31.11% 0.1424 0.0990 30.48% 0.5686 0.4166 26.73%
Qwen2.5-32B-Instruct 0.4922 0.3244 34.09% 0.4926 0.3462 29.72% 0.1376 0.1004 27.03% 0.5830 0.4668 19.93%
Qwen2.5-3B-Instruct 0.4146 0.2792 32.66% 0.3970 0.2632 33.70% 0.1274 0.0856 32.81% 0.5134 0.3940 23.26%
Qwen2.5-72B-Instruct 0.4968 0.3278 34.02% 0.5066 0.3558 29.77% 0.1316 0.0958 27.20% 0.5956 0.4784 19.68%
Qwen2.5-7B-Instruct 0.4544 0.3082 32.17% 0.4474 0.2988 33.21% 0.1406 0.0956 32.01% 0.5630 0.4320 23.27%
Contriever, Typos, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.4942 0.3894 21.21% 0.4808 0.4120 14.31% 0.1270 0.1062 16.38% 0.5914 0.5232 11.53%
gemma-2-2b-it 0.4628 0.3458 25.28% 0.4090 0.3196 21.86% 0.1282 0.1036 19.19% 0.5568 0.4794 13.90%
gemma-2-9b-it 0.4746 0.3660 22.88% 0.4512 0.3684 18.35% 0.1304 0.1078 17.33% 0.5672 0.4974 12.31%
Llama-3.1-70B-Instruct 0.5064 0.3886 23.26% 0.4866 0.3830 21.29% 0.1420 0.1164 18.03% 0.5818 0.4916 15.50%
Llama-3.1-8B-Instruct 0.4708 0.3460 26.51% 0.4622 0.3650 21.03% 0.1424 0.1148 19.38% 0.5686 0.4754 16.39%
Qwen2.5-32B-Instruct 0.4922 0.3654 25.76% 0.4926 0.3992 18.96% 0.1376 0.1146 16.72% 0.5830 0.4920 15.61%
Qwen2.5-3B-Instruct 0.4146 0.3038 26.72% 0.3970 0.2868 27.76% 0.1274 0.0952 25.27% 0.5134 0.4196 18.27%
Qwen2.5-72B-Instruct 0.4976 0.3796 23.71% 0.5066 0.4336 14.41% 0.1316 0.1104 16.11% 0.5936 0.5198 12.43%
Qwen2.5-7B-Instruct 0.4544 0.3348 26.32% 0.4474 0.3434 23.25% 0.1406 0.1124 20.06% 0.5630 0.4866 13.57%
Table 9: RAG experiment results with Contriever as retrieval model on AM scores.
Contriever, Formality, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.1454 0.0151 89.64% 0.0550 0.0205 62.67% 0.0196 0.0058 70.41% 0.1914 0.0330 82.76%
gemma-2-9b-it 0.0214 0.0066 69.16% 0.0196 0.0093 52.72% 0.0340 0.0089 73.92% 0.0888 0.0341 61.64%
gemma-2-27b-it 0.0044 0.0122 177.27% 0.0188 0.0147 21.63% 0.0058 0.0037 36.78% 0.0292 0.0505 72.83%
Llama-3.1-8B-Instruct 0.3282 0.1348 58.93% 0.1482 0.0525 64.60% 0.0216 0.0055 74.69% 0.2896 0.1429 50.67%
Llama-3.1-70B-Instruct 0.3432 0.0884 74.24% 0.1234 0.0369 70.12% 0.0244 0.0069 71.86% 0.2996 0.1177 60.70%
Qwen2.5-3B-Instruct 0.3906 0.1705 56.34% 0.1810 0.0842 53.48% 0.0186 0.0058 68.82% 0.3404 0.2079 38.93%
Qwen2.5-7B-Instruct 0.4242 0.1847 56.45% 0.1892 0.1138 39.85% 0.0352 0.0148 57.95% 0.2862 0.1521 46.87%
Qwen2.5-32B-Instruct 0.3736 0.1257 66.35% 0.1086 0.0609 43.95% 0.0236 0.0105 55.37% 0.3580 0.1762 50.78%
Qwen2.5-72B-Instruct 0.4656 0.1781 61.76% 0.1490 0.0733 50.83% 0.0204 0.0094 53.92% 0.3890 0.1992 48.79%
Contriever, Politeness, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0034 0.0152 347.06% 0.0148 0.0207 40.09% 0.0060 0.0054 10.00% 0.0188 0.0259 37.59%
gemma-2-2b-it 0.0970 0.0410 57.73% 0.0450 0.0424 5.78% 0.0200 0.0082 59.00% 0.1752 0.0358 79.57%
gemma-2-9b-it 0.0170 0.0068 60.00% 0.0130 0.0105 19.49% 0.0512 0.0153 70.18% 0.0912 0.0278 69.52%
Llama-3.1-70B-Instruct 0.1906 0.0875 54.11% 0.1064 0.0436 59.02% 0.0236 0.0082 65.25% 0.2330 0.1326 43.09%
Llama-3.1-8B-Instruct 0.1998 0.1657 17.05% 0.1024 0.0673 34.31% 0.0272 0.0077 71.57% 0.2840 0.2535 10.73%
Qwen2.5-32B-Instruct 0.2020 0.1683 16.70% 0.0546 0.0851 55.80% 0.0192 0.0155 19.44% 0.3344 0.3055 8.65%
Qwen2.5-3B-Instruct 0.2694 0.2473 8.22% 0.1920 0.1695 11.70% 0.0236 0.0159 32.77% 0.3338 0.3331 0.22%
Qwen2.5-72B-Instruct 0.3270 0.2855 12.68% 0.1030 0.1504 46.02% 0.0242 0.0187 22.59% 0.3912 0.3877 0.90%
Qwen2.5-7B-Instruct 0.2736 0.2719 0.61% 0.1700 0.1984 16.71% 0.0384 0.0261 32.12% 0.3160 0.3382 7.03%
Contriever, Readability, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0008 0.0010 25.00% 0.0056 0.0054 3.57% 0.0042 0.0035 17.46% 0.0062 0.0024 61.29%
gemma-2-2b-it 0.0928 0.0661 28.81% 0.0282 0.0414 46.81% 0.0134 0.0205 52.74% 0.0674 0.0692 2.67%
gemma-2-9b-it 0.0114 0.0379 232.16% 0.0072 0.0223 210.19% 0.0454 0.0521 14.68% 0.0262 0.0545 107.89%
Llama-3.1-70B-Instruct 0.2596 0.0733 71.78% 0.0962 0.0543 43.59% 0.0176 0.0071 59.85% 0.1228 0.0776 36.81%
Llama-3.1-8B-Instruct 0.2438 0.1507 38.17% 0.1024 0.0709 30.79% 0.0168 0.0071 57.94% 0.1412 0.1187 15.96%
Qwen2.5-32B-Instruct 0.2188 0.1020 53.38% 0.0472 0.0719 52.26% 0.0126 0.0137 8.99% 0.1516 0.1405 7.30%
Qwen2.5-3B-Instruct 0.2910 0.1398 51.96% 0.1902 0.1187 37.57% 0.0214 0.0152 28.97% 0.1694 0.1279 24.52%
Qwen2.5-72B-Instruct 0.3788 0.2529 33.23% 0.1002 0.1301 29.87% 0.0168 0.0151 10.32% 0.2200 0.2015 8.39%
Qwen2.5-7B-Instruct 0.3244 0.1885 41.90% 0.1476 0.1629 10.39% 0.0348 0.0441 26.82% 0.1354 0.1303 3.79%
Contriever, Round-trip Translation, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0004 0.0012 200.00% 0.0144 0.0164 13.89% 0.0046 0.0058 26.09% 0.0104 0.0116 11.54%
gemma-2-2b-it 0.0836 0.0630 24.64% 0.0474 0.0288 39.24% 0.0126 0.0102 19.05% 0.0738 0.0582 21.14%
gemma-2-9b-it 0.0078 0.0062 20.51% 0.0176 0.0154 12.50% 0.0380 0.0294 22.63% 0.0552 0.0358 35.14%
Llama-3.1-70B-Instruct 0.2350 0.1274 45.79% 0.0324 0.0876 170.37% 0.0172 0.0158 8.14% 0.1742 0.1222 29.85%
Llama-3.1-8B-Instruct 0.2500 0.1400 44.00% 0.1614 0.1072 33.58% 0.0306 0.0150 50.98% 0.2052 0.1302 36.55%
Qwen2.5-32B-Instruct 0.1360 0.0658 51.62% 0.0800 0.0518 35.25% 0.0100 0.0074 26.00% 0.1734 0.1328 23.41%
Qwen2.5-3B-Instruct 0.2104 0.1042 50.48% 0.1652 0.0850 48.55% 0.0202 0.0158 21.78% 0.2234 0.1606 28.11%
Qwen2.5-72B-Instruct 0.3178 0.1694 46.70% 0.1460 0.0844 42.19% 0.0154 0.0116 24.68% 0.2850 0.2220 22.11%
Qwen2.5-7B-Instruct 0.2438 0.1452 40.44% 0.1560 0.1174 24.74% 0.0334 0.0272 18.56% 0.1536 0.1264 17.71%
Contriever, Typos, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0004 0.0010 150.00% 0.0144 0.0124 13.89% 0.0046 0.0024 47.83% 0.0104 0.0154 48.08%
gemma-2-2b-it 0.0836 0.0588 29.67% 0.0474 0.0338 28.69% 0.0126 0.0114 9.52% 0.0738 0.0556 24.66%
gemma-2-9b-it 0.0078 0.0084 7.69% 0.0176 0.0134 23.86% 0.0380 0.0300 21.05% 0.0552 0.0548 0.72%
Llama-3.1-70B-Instruct 0.2350 0.1466 37.62% 0.1302 0.0944 27.50% 0.0172 0.0118 31.40% 0.1742 0.1302 25.26%
Llama-3.1-8B-Instruct 0.2500 0.1676 32.96% 0.1614 0.1106 31.47% 0.0306 0.0186 39.22% 0.2052 0.1742 15.11%
Qwen2.5-32B-Instruct 0.1360 0.0848 37.65% 0.0800 0.0548 31.50% 0.0100 0.0084 16.00% 0.1734 0.1348 22.26%
Qwen2.5-3B-Instruct 0.2104 0.1314 37.55% 0.1652 0.1064 35.59% 0.0202 0.0156 22.77% 0.2234 0.1720 23.01%
Qwen2.5-72B-Instruct 0.3160 0.2100 33.54% 0.1460 0.0962 34.11% 0.0154 0.0110 28.57% 0.2784 0.2264 18.68%
Qwen2.5-7B-Instruct 0.2438 0.1512 37.98% 0.1560 0.1028 34.10% 0.0334 0.0268 19.76% 0.1536 0.1144 25.52%
Table 10: RAG experiment results with Contriever as retrieval model on EM scores.
Contriever, Formality, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.3457 0.2167 37.31% 0.2285 0.1737 23.99% 0.2474 0.1933 21.86% 0.3694 0.2298 37.79%
gemma-2-9b-it 0.2809 0.2274 19.06% 0.2142 0.1911 10.75% 0.2614 0.1988 23.97% 0.2996 0.2436 18.70%
gemma-2-27b-it 0.2756 0.2364 14.24% 0.2193 0.1972 10.10% 0.2432 0.2016 17.10% 0.2573 0.2569 0.12%
Llama-3.1-8B-Instruct 0.4601 0.2842 38.22% 0.3019 0.1921 36.37% 0.2478 0.1910 22.93% 0.4349 0.2964 31.85%
Llama-3.1-70B-Instruct 0.4676 0.2610 44.17% 0.2861 0.1858 35.06% 0.2502 0.1884 24.71% 0.4406 0.2836 35.65%
Qwen2.5-3B-Instruct 0.4801 0.2843 40.77% 0.3212 0.2015 37.27% 0.2401 0.1764 26.55% 0.4530 0.3201 29.34%
Qwen2.5-7B-Instruct 0.5191 0.3203 38.30% 0.3464 0.2469 28.73% 0.2601 0.1970 24.25% 0.4234 0.3041 28.18%
Qwen2.5-32B-Instruct 0.4835 0.2844 41.18% 0.2789 0.2072 25.71% 0.2541 0.2003 21.15% 0.4824 0.3295 31.69%
Qwen2.5-72B-Instruct 0.5521 0.3208 41.90% 0.3159 0.2203 30.26% 0.2525 0.2025 19.80% 0.5197 0.3480 33.05%
Contriever, Politeness, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2376 0.2257 5.00% 0.2153 0.2158 0.22% 0.2418 0.2230 7.79% 0.2395 0.2443 1.99%
gemma-2-2b-it 0.2789 0.2273 18.52% 0.2135 0.2042 4.36% 0.2452 0.2171 11.45% 0.3630 0.2491 31.39%
gemma-2-9b-it 0.2354 0.2031 13.72% 0.2051 0.1961 4.36% 0.2722 0.2271 16.58% 0.2964 0.2465 16.85%
Llama-3.1-70B-Instruct 0.3254 0.2443 24.93% 0.2698 0.2022 25.05% 0.2474 0.2135 13.72% 0.3954 0.3075 22.24%
Llama-3.1-8B-Instruct 0.3312 0.2908 12.19% 0.2593 0.2179 15.98% 0.2504 0.2146 14.29% 0.4461 0.4117 7.70%
Qwen2.5-32B-Instruct 0.3381 0.2963 12.38% 0.2361 0.2474 4.82% 0.2465 0.2241 9.07% 0.4744 0.4482 5.51%
Qwen2.5-3B-Instruct 0.3592 0.3223 10.28% 0.3236 0.2917 9.86% 0.2416 0.2129 11.88% 0.4636 0.4574 1.33%
Qwen2.5-72B-Instruct 0.4211 0.3810 9.51% 0.2784 0.3069 10.26% 0.2503 0.2254 9.96% 0.5414 0.5366 0.89%
Qwen2.5-7B-Instruct 0.3806 0.3632 4.59% 0.3241 0.3359 3.65% 0.2606 0.2300 11.73% 0.4635 0.4804 3.65%
Contriever, Readability, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2390 0.1688 29.36% 0.2089 0.1806 13.55% 0.2299 0.1939 15.65% 0.2213 0.1890 14.59%
gemma-2-2b-it 0.2754 0.1968 28.55% 0.1968 0.1752 10.99% 0.2304 0.1959 14.97% 0.2739 0.2429 11.34%
gemma-2-9b-it 0.2385 0.1894 20.58% 0.2010 0.1846 8.15% 0.2569 0.2245 12.64% 0.2369 0.2295 3.11%
Llama-3.1-70B-Instruct 0.3804 0.2019 46.92% 0.2591 0.1865 28.03% 0.2331 0.1837 21.19% 0.2926 0.2267 22.52%
Llama-3.1-8B-Instruct 0.3682 0.2547 30.83% 0.2568 0.1963 23.54% 0.2304 0.1822 20.92% 0.3168 0.2756 12.99%
Qwen2.5-32B-Instruct 0.3543 0.2240 36.78% 0.2271 0.2059 9.33% 0.2316 0.1958 15.46% 0.3115 0.2857 8.29%
Qwen2.5-3B-Instruct 0.3817 0.2532 33.66% 0.3122 0.2464 21.08% 0.2309 0.1848 19.97% 0.3055 0.2682 12.20%
Qwen2.5-72B-Instruct 0.4681 0.3387 27.64% 0.2716 0.2601 4.24% 0.2376 0.2015 15.16% 0.3820 0.3439 9.96%
Qwen2.5-7B-Instruct 0.4173 0.2902 30.45% 0.2957 0.2902 1.84% 0.2493 0.2212 11.25% 0.3021 0.2848 5.73%
Contriever, Round-trip Translation, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2156 0.1322 38.67% 0.2102 0.1594 24.20% 0.2151 0.1807 15.96% 0.2206 0.1888 14.41%
gemma-2-2b-it 0.2574 0.1672 35.03% 0.2119 0.1448 31.65% 0.2182 0.1789 17.99% 0.2719 0.2216 18.51%
gemma-2-9b-it 0.2116 0.1302 38.45% 0.2041 0.1539 24.60% 0.2389 0.1963 17.84% 0.2550 0.2059 19.25%
Llama-3.1-70B-Instruct 0.3436 0.2038 40.69% 0.2321 0.2024 12.80% 0.2223 0.1857 16.44% 0.3382 0.2582 23.65%
Llama-3.1-8B-Instruct 0.3534 0.2143 39.36% 0.3134 0.2214 29.37% 0.2318 0.1816 21.67% 0.3654 0.2492 31.80%
Qwen2.5-32B-Instruct 0.2837 0.1646 41.99% 0.2496 0.1733 30.57% 0.2199 0.1841 16.27% 0.3320 0.2639 20.52%
Qwen2.5-3B-Instruct 0.3061 0.1789 41.55% 0.2998 0.1825 39.11% 0.2162 0.1720 20.44% 0.3651 0.2859 21.70%
Qwen2.5-72B-Instruct 0.4089 0.2375 41.91% 0.3098 0.2054 33.71% 0.2248 0.1896 15.65% 0.4348 0.3451 20.62%
Qwen2.5-7B-Instruct 0.3480 0.2204 36.66% 0.3090 0.2270 26.53% 0.2389 0.1952 18.27% 0.3230 0.2672 17.27%
Contriever, Typos, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2156 0.1658 23.08% 0.2102 0.1852 11.90% 0.2151 0.1951 9.28% 0.2206 0.2006 9.09%
gemma-2-2b-it 0.2574 0.1878 27.05% 0.2119 0.1701 19.74% 0.2182 0.1975 9.50% 0.2719 0.2308 15.12%
gemma-2-9b-it 0.2116 0.1604 24.20% 0.2041 0.1719 15.76% 0.2389 0.2114 11.52% 0.2550 0.2295 10.00%
Llama-3.1-70B-Instruct 0.3436 0.2450 28.71% 0.2885 0.2269 21.35% 0.2223 0.1979 10.95% 0.3382 0.2710 19.87%
Llama-3.1-8B-Instruct 0.3534 0.2499 29.27% 0.3134 0.2400 23.44% 0.2318 0.2031 12.38% 0.3654 0.3107 14.97%
Qwen2.5-32B-Instruct 0.2837 0.1982 30.13% 0.2496 0.1992 20.18% 0.2199 0.1979 9.99% 0.3320 0.2723 17.98%
Qwen2.5-3B-Instruct 0.3061 0.2093 31.61% 0.2998 0.2132 28.89% 0.2162 0.1874 13.31% 0.3651 0.2920 20.04%
Qwen2.5-72B-Instruct 0.4067 0.2903 28.62% 0.3098 0.2488 19.71% 0.2248 0.2083 7.36% 0.4299 0.3632 15.51%
Qwen2.5-7B-Instruct 0.3480 0.2378 31.67% 0.3090 0.2324 24.79% 0.2389 0.2107 11.82% 0.3230 0.2613 19.11%
Table 11: RAG experiment results with Contriever as retrieval model on F1 scores.
ModernBERT, Formality, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.6126 0.5291 13.64% 0.4608 0.4035 12.44% 0.1690 0.1296 23.31% 0.6012 0.5551 7.66%
gemma-2-9b-it 0.6360 0.5618 11.67% 0.4984 0.4488 9.95% 0.1724 0.1385 19.68% 0.6178 0.5631 8.85%
gemma-2-27b-it 0.6508 0.5713 12.21% 0.5230 0.4711 9.92% 0.1676 0.1355 19.13% 0.6264 0.5813 7.19%
Llama-3.1-8B-Instruct 0.6488 0.5695 12.23% 0.5228 0.4679 10.51% 0.1982 0.1633 17.59% 0.6158 0.5717 7.16%
Llama-3.1-70B-Instruct 0.6722 0.5950 11.48% 0.5378 0.4728 12.09% 0.1986 0.1609 19.00% 0.6350 0.5893 7.19%
Qwen2.5-3B-Instruct 0.5806 0.5169 10.97% 0.4534 0.4018 11.38% 0.1834 0.1451 20.87% 0.5618 0.5196 7.51%
Qwen2.5-7B-Instruct 0.6290 0.5473 12.99% 0.5010 0.4436 11.46% 0.1848 0.1481 19.84% 0.6146 0.5649 8.08%
Qwen2.5-32B-Instruct 0.6512 0.5773 11.34% 0.5274 0.4860 7.85% 0.1766 0.1505 14.80% 0.6206 0.5819 6.24%
Qwen2.5-72B-Instruct 0.6576 0.5874 10.68% 0.5468 0.4970 9.11% 0.1776 0.1408 20.72% 0.6312 0.5949 5.75%
ModernBERT, Politeness, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5632 0.5185 7.93% 0.5132 0.4959 3.38% 0.1958 0.1838 6.13% 0.6138 0.6017 1.97%
gemma-2-2b-it 0.5220 0.4783 8.37% 0.4394 0.4200 4.42% 0.2020 0.1795 11.16% 0.5938 0.5762 2.96%
gemma-2-9b-it 0.5394 0.4933 8.54% 0.4830 0.4656 3.60% 0.2044 0.1877 8.19% 0.6054 0.5872 3.01%
Llama-3.1-70B-Instruct 0.5734 0.5333 7.00% 0.5328 0.4993 6.29% 0.2306 0.2078 9.89% 0.6124 0.5903 3.61%
Llama-3.1-8B-Instruct 0.5498 0.5081 7.59% 0.5080 0.4904 3.46% 0.2302 0.2074 9.90% 0.5912 0.5705 3.50%
Qwen2.5-32B-Instruct 0.5602 0.5117 8.65% 0.5154 0.4909 4.76% 0.2062 0.1925 6.66% 0.6020 0.5681 5.64%
Qwen2.5-3B-Instruct 0.4712 0.4168 11.54% 0.4250 0.3999 5.91% 0.2130 0.1868 12.30% 0.5506 0.5080 7.74%
Qwen2.5-72B-Instruct 0.5626 0.5212 7.36% 0.5398 0.5104 5.45% 0.1970 0.1643 16.62% 0.6182 0.5943 3.86%
Qwen2.5-7B-Instruct 0.5360 0.4783 10.76% 0.4782 0.4541 5.05% 0.2128 0.1977 7.08% 0.5962 0.5683 4.69%
ModernBERT, Readability, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5592 0.4689 16.15% 0.5220 0.4629 11.33% 0.1968 0.1804 8.33% 0.5480 0.4785 12.68%
gemma-2-2b-it 0.5180 0.4107 20.71% 0.4460 0.3710 16.82% 0.2038 0.1683 17.44% 0.5088 0.4252 16.43%
gemma-2-9b-it 0.5394 0.4471 17.12% 0.4926 0.4239 13.95% 0.1970 0.1751 11.13% 0.5356 0.4539 15.25%
Llama-3.1-70B-Instruct 0.5768 0.4749 17.66% 0.5436 0.4579 15.77% 0.2226 0.1930 13.30% 0.5522 0.4633 16.09%
Llama-3.1-8B-Instruct 0.5546 0.4588 17.27% 0.5180 0.4399 15.07% 0.2210 0.1873 15.23% 0.5194 0.4503 13.30%
Qwen2.5-32B-Instruct 0.5550 0.4607 17.00% 0.5234 0.4501 14.01% 0.2800 0.1901 32.10% 0.5310 0.4333 18.41%
Qwen2.5-3B-Instruct 0.4742 0.3682 22.35% 0.4232 0.3443 18.65% 0.2070 0.1795 13.27% 0.4560 0.3459 24.14%
Qwen2.5-72B-Instruct 0.5620 0.4707 16.25% 0.5444 0.4703 13.62% 0.2042 0.1871 8.36% 0.5352 0.4441 17.03%
Qwen2.5-7B-Instruct 0.5286 0.4255 19.50% 0.4832 0.4003 17.15% 0.2118 0.1873 11.55% 0.5104 0.4065 20.35%
ModernBERT, Round-trip Translation, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5304 0.3564 32.81% 0.5052 0.3838 24.03% 0.1530 0.1106 27.71% 0.5836 0.4626 20.73%
gemma-2-2b-it 0.4876 0.3280 32.73% 0.4272 0.3072 28.09% 0.1590 0.1116 29.81% 0.5384 0.4174 22.47%
gemma-2-9b-it 0.5092 0.3366 33.90% 0.4760 0.3594 24.50% 0.1604 0.1192 25.69% 0.5602 0.4318 22.92%
Llama-3.1-70B-Instruct 0.5388 0.3486 35.30% 0.5182 0.3864 25.43% 0.1748 0.1314 24.83% 0.5728 0.4490 21.61%
Llama-3.1-8B-Instruct 0.5030 0.3256 35.27% 0.4954 0.3594 27.45% 0.1860 0.1258 32.37% 0.5554 0.4162 25.06%
Qwen2.5-32B-Instruct 0.5230 0.3446 34.11% 0.5170 0.3888 24.80% 0.1642 0.1204 26.67% 0.5678 0.4488 20.96%
Qwen2.5-3B-Instruct 0.4366 0.2930 32.89% 0.4372 0.3028 30.74% 0.1674 0.1136 32.14% 0.4908 0.3704 24.53%
Qwen2.5-72B-Instruct 0.5314 0.3524 33.68% 0.5272 0.3992 24.28% 0.1558 0.1130 27.47% 0.5798 0.4664 19.56%
Qwen2.5-7B-Instruct 0.4950 0.3264 34.06% 0.4854 0.3414 29.67% 0.1656 0.1254 24.28% 0.5522 0.4320 21.77%
ModernBERT, Typos, AM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.5304 0.4338 18.21% 0.5052 0.4708 6.81% 0.1530 0.1472 3.79% 0.5836 0.5212 10.69%
gemma-2-2b-it 0.4876 0.3888 20.26% 0.4272 0.3826 10.44% 0.1590 0.1428 10.19% 0.5384 0.4826 10.36%
gemma-2-9b-it 0.5092 0.4094 19.60% 0.4760 0.4370 8.19% 0.1604 0.1492 6.98% 0.5602 0.5020 10.39%
Llama-3.1-70B-Instruct 0.5388 0.4338 19.49% 0.5182 0.4726 8.80% 0.1748 0.1654 5.38% 0.5728 0.5000 12.71%
Llama-3.1-8B-Instruct 0.5030 0.3926 21.95% 0.4954 0.4432 10.54% 0.1860 0.1630 12.37% 0.5554 0.4858 12.53%
Qwen2.5-32B-Instruct 0.5230 0.4200 19.69% 0.5170 0.4690 9.28% 0.1642 0.1522 7.31% 0.5678 0.5018 11.62%
Qwen2.5-3B-Instruct 0.4366 0.3328 23.77% 0.4372 0.3688 15.65% 0.1674 0.1468 12.31% 0.4908 0.4260 13.20%
Qwen2.5-72B-Instruct 0.5314 0.4290 19.27% 0.5272 0.4864 7.74% 0.1558 0.1430 8.22% 0.5798 0.5198 10.35%
Qwen2.5-7B-Instruct 0.4950 0.3828 22.67% 0.4854 0.4214 13.19% 0.1656 0.1546 6.64% 0.5522 0.4860 11.99%
Table 12: RAG experiment results with ModernBERT as retrieval model on AM scores.
ModernBERT, Formality, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.1554 0.0167 89.23% 0.0558 0.0204 63.44% 0.0220 0.0066 70.00% 0.1964 0.0299 84.76%
gemma-2-9b-it 0.0282 0.0072 74.47% 0.0210 0.0091 56.83% 0.0358 0.0096 73.18% 0.0992 0.0356 64.11%
gemma-2-27b-it 0.0052 0.0105 101.28% 0.0202 0.0141 30.36% 0.0096 0.0039 59.03% 0.0320 0.0496 55.00%
Llama-3.1-8B-Instruct 0.3500 0.1525 56.42% 0.1642 0.0551 66.42% 0.0286 0.0115 59.67% 0.2886 0.1429 50.47%
Llama-3.1-70B-Instruct 0.3748 0.1008 73.11% 0.1342 0.0375 72.03% 0.0324 0.0123 61.93% 0.3088 0.1203 61.03%
Qwen2.5-3B-Instruct 0.4016 0.1943 51.63% 0.1878 0.0921 50.98% 0.0248 0.0095 61.56% 0.3384 0.2162 36.11%
Qwen2.5-7B-Instruct 0.4554 0.2097 53.96% 0.2116 0.1267 40.11% 0.0460 0.0224 51.30% 0.2936 0.1527 48.00%
Qwen2.5-32B-Instruct 0.3986 0.1289 67.67% 0.1096 0.0643 41.36% 0.0260 0.0139 46.67% 0.3562 0.1733 51.34%
Qwen2.5-72B-Instruct 0.4924 0.2003 59.33% 0.1612 0.0794 50.74% 0.0254 0.0129 49.34% 0.3920 0.1996 49.08%
ModernBERT, Politeness, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0038 0.0124 226.32% 0.0132 0.0233 76.26% 0.0080 0.0056 30.00% 0.0166 0.0294 77.11%
gemma-2-2b-it 0.1046 0.0437 58.19% 0.0476 0.0434 8.82% 0.0218 0.0101 53.52% 0.1774 0.0355 79.97%
gemma-2-9b-it 0.0218 0.0077 64.83% 0.0122 0.0111 8.74% 0.0514 0.0151 70.69% 0.0932 0.0333 64.31%
Llama-3.1-70B-Instruct 0.2240 0.0986 55.98% 0.1106 0.0487 55.94% 0.0274 0.0129 53.04% 0.2366 0.1333 43.65%
Llama-3.1-8B-Instruct 0.2158 0.1763 18.32% 0.1006 0.0710 29.42% 0.0312 0.0135 56.84% 0.2864 0.2413 15.74%
Qwen2.5-32B-Instruct 0.2198 0.1857 15.50% 0.0608 0.0904 48.68% 0.0192 0.0165 13.89% 0.3330 0.2845 14.55%
Qwen2.5-3B-Instruct 0.2916 0.2636 9.60% 0.2076 0.1827 11.98% 0.0298 0.0213 28.41% 0.3230 0.3073 4.85%
Qwen2.5-72B-Instruct 0.3540 0.3109 12.18% 0.1166 0.1603 37.51% 0.0276 0.0224 18.84% 0.3928 0.3777 3.84%
Qwen2.5-7B-Instruct 0.3170 0.2979 6.01% 0.1968 0.2121 7.76% 0.0474 0.0355 25.18% 0.3184 0.3227 1.36%
ModernBERT, Readability, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0010 0.0013 26.67% 0.0068 0.0049 27.45% 0.0054 0.0032 40.74% 0.0046 0.0027 42.03%
gemma-2-2b-it 0.0852 0.0649 23.87% 0.0316 0.0445 40.93% 0.0166 0.0204 22.89% 0.0660 0.0655 0.71%
gemma-2-9b-it 0.0156 0.0433 177.35% 0.0086 0.0254 195.35% 0.0460 0.0554 20.43% 0.0316 0.0561 77.64%
Llama-3.1-70B-Instruct 0.2854 0.0745 73.91% 0.1008 0.0564 44.05% 0.0208 0.0089 57.05% 0.1214 0.0753 38.00%
Llama-3.1-8B-Instruct 0.2542 0.1536 39.58% 0.1096 0.0707 35.52% 0.0230 0.0092 60.00% 0.1346 0.1112 17.38%
Qwen2.5-32B-Instruct 0.2272 0.1111 51.09% 0.0556 0.0771 38.61% 0.0100 0.0164 64.00% 0.1442 0.1272 11.79%
Qwen2.5-3B-Instruct 0.3164 0.1418 55.18% 0.2158 0.1288 40.32% 0.0298 0.0206 30.87% 0.1588 0.1095 31.02%
Qwen2.5-72B-Instruct 0.4048 0.2793 30.99% 0.1112 0.1393 25.30% 0.0220 0.0211 4.24% 0.2174 0.1915 11.93%
Qwen2.5-7B-Instruct 0.3544 0.2018 43.06% 0.1648 0.1814 10.07% 0.0408 0.0569 39.54% 0.1336 0.1189 10.98%
ModernBERT, Round-trip Translation, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0006 0.0012 100.00% 0.0172 0.0182 5.81% 0.0068 0.0050 26.47% 0.0078 0.0086 10.26%
gemma-2-2b-it 0.0754 0.0616 18.30% 0.0474 0.0328 30.80% 0.0154 0.0124 19.48% 0.0736 0.0500 32.07%
gemma-2-9b-it 0.0098 0.0118 20.41% 0.0186 0.0178 4.30% 0.0416 0.0298 28.37% 0.0574 0.0378 34.15%
Llama-3.1-70B-Instruct 0.2532 0.1396 44.87% 0.1416 0.1120 20.90% 0.0248 0.0202 18.55% 0.1774 0.1250 29.54%
Llama-3.1-8B-Instruct 0.2660 0.1550 41.73% 0.1792 0.1284 28.35% 0.0402 0.0232 42.29% 0.1968 0.1274 35.26%
Qwen2.5-32B-Instruct 0.1398 0.0706 49.50% 0.0820 0.0572 30.24% 0.0126 0.0102 19.05% 0.1686 0.1160 31.20%
Qwen2.5-3B-Instruct 0.2282 0.1198 47.50% 0.1818 0.1096 39.71% 0.0276 0.0182 34.06% 0.2036 0.1344 33.99%
Qwen2.5-72B-Instruct 0.3356 0.1938 42.25% 0.1618 0.0984 39.18% 0.0208 0.0144 30.77% 0.2700 0.2146 20.52%
Qwen2.5-7B-Instruct 0.2756 0.1582 42.60% 0.1676 0.1390 17.06% 0.0418 0.0306 26.79% 0.1522 0.1196 21.42%
ModernBERT, Typos, EM Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.0006 0.0008 33.33% 0.0172 0.0094 45.35% 0.0068 0.0052 23.53% 0.0078 0.0088 12.82%
gemma-2-2b-it 0.0754 0.0642 14.85% 0.0474 0.0408 13.92% 0.0154 0.0156 1.30% 0.0736 0.0542 26.36%
gemma-2-9b-it 0.0098 0.0096 2.04% 0.0186 0.0132 29.03% 0.0416 0.0342 17.79% 0.0574 0.0532 7.32%
Llama-3.1-70B-Instruct 0.2532 0.1634 35.47% 0.1416 0.1006 28.95% 0.0248 0.0196 20.97% 0.1774 0.1358 23.45%
Llama-3.1-8B-Instruct 0.2660 0.1850 30.45% 0.1792 0.1358 24.22% 0.0402 0.0322 19.90% 0.1968 0.1826 7.22%
Qwen2.5-32B-Instruct 0.1398 0.0940 32.76% 0.0820 0.0642 21.71% 0.0126 0.0112 11.11% 0.1686 0.1344 20.28%
Qwen2.5-3B-Instruct 0.2282 0.1568 31.29% 0.1818 0.1448 20.35% 0.0276 0.0268 2.90% 0.2036 0.1758 13.65%
Qwen2.5-72B-Instruct 0.3356 0.2424 27.77% 0.1618 0.1144 29.30% 0.0208 0.0160 23.08% 0.2700 0.2288 15.26%
Qwen2.5-7B-Instruct 0.2756 0.1832 33.53% 0.1676 0.1290 23.03% 0.0418 0.0342 18.18% 0.1522 0.1180 22.47%
Table 13: RAG experiment results with ModernBERT as retrieval model on EM scores.
ModernBERT, Formality, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-2b-it 0.3666 0.2525 31.14% 0.2346 0.1833 21.88% 0.2800 0.2276 18.73% 0.3696 0.2327 37.05%
gemma-2-9b-it 0.3018 0.2642 12.48% 0.2217 0.2007 9.46% 0.2927 0.2429 17.01% 0.3082 0.2511 18.55%
gemma-2-27b-it 0.2965 0.2718 8.33% 0.2273 0.2053 9.67% 0.2746 0.2416 12.03% 0.2574 0.2625 1.96%
Llama-3.1-8B-Instruct 0.4871 0.3223 33.84% 0.3205 0.2005 37.43% 0.2880 0.2396 16.81% 0.4291 0.2985 30.45%
Llama-3.1-70B-Instruct 0.4998 0.2968 40.63% 0.2993 0.1972 34.10% 0.2931 0.2467 15.83% 0.4452 0.2891 35.07%
Qwen2.5-3B-Instruct 0.4911 0.3256 33.71% 0.3366 0.2176 35.34% 0.2789 0.2208 20.84% 0.4476 0.3300 26.27%
Qwen2.5-7B-Instruct 0.5522 0.3675 33.45% 0.3655 0.2679 26.72% 0.3030 0.2440 19.48% 0.4276 0.3098 27.55%
Qwen2.5-32B-Instruct 0.5138 0.3227 37.20% 0.2845 0.2176 23.52% 0.2814 0.2356 16.26% 0.4781 0.3325 30.46%
Qwen2.5-72B-Instruct 0.5824 0.3645 37.42% 0.3298 0.2298 30.33% 0.2848 0.2374 16.64% 0.5187 0.3516 32.22%
ModernBERT, Politeness, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2609 0.2398 8.08% 0.2241 0.2211 1.33% 0.2740 0.2577 5.96% 0.2376 0.2450 3.09%
gemma-2-2b-it 0.3058 0.2442 20.16% 0.2256 0.2057 8.84% 0.2790 0.2551 8.58% 0.3653 0.2425 33.61%
gemma-2-9b-it 0.2625 0.2182 16.86% 0.2150 0.2006 6.70% 0.3068 0.2649 13.66% 0.3005 0.2472 17.74%
Llama-3.1-70B-Instruct 0.3630 0.2654 26.88% 0.2822 0.2108 25.30% 0.2901 0.2616 9.82% 0.4009 0.3021 24.65%
Llama-3.1-8B-Instruct 0.3572 0.3119 12.67% 0.2635 0.2222 15.69% 0.2931 0.2596 11.43% 0.4443 0.3962 10.83%
Qwen2.5-32B-Instruct 0.3720 0.3239 12.95% 0.2478 0.2536 2.34% 0.2747 0.2530 7.90% 0.4738 0.4264 10.01%
Qwen2.5-3B-Instruct 0.3851 0.3427 11.02% 0.3434 0.3042 11.40% 0.2806 0.2531 9.81% 0.4545 0.4304 5.31%
Qwen2.5-72B-Instruct 0.4562 0.4112 9.85% 0.2949 0.3168 7.44% 0.2827 0.2571 9.09% 0.5433 0.5237 3.61%
Qwen2.5-7B-Instruct 0.4284 0.3926 8.37% 0.3486 0.3507 0.59% 0.3010 0.2737 9.06% 0.4653 0.4628 0.53%
ModernBERT, Readability, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2571 0.1885 26.69% 0.2185 0.1850 15.36% 0.2555 0.2220 13.10% 0.2156 0.1820 15.59%
gemma-2-2b-it 0.2911 0.2113 27.41% 0.2097 0.1815 13.45% 0.2595 0.2220 14.44% 0.2660 0.2256 15.20%
gemma-2-9b-it 0.2613 0.2084 20.24% 0.2122 0.1946 8.32% 0.2868 0.2596 9.49% 0.2378 0.2210 7.09%
Llama-3.1-70B-Instruct 0.4085 0.2158 47.16% 0.2710 0.1934 28.63% 0.2673 0.2209 17.38% 0.2904 0.2170 25.28%
Llama-3.1-8B-Instruct 0.3897 0.2701 30.69% 0.2701 0.1983 26.57% 0.2662 0.2148 19.30% 0.3075 0.2561 16.71%
Qwen2.5-32B-Instruct 0.3757 0.2450 34.78% 0.2402 0.2137 11.02% 0.2595 0.2248 13.36% 0.3007 0.2617 12.97%
Qwen2.5-3B-Instruct 0.4077 0.2696 33.89% 0.3384 0.2608 22.93% 0.2664 0.2218 16.74% 0.2895 0.2402 17.02%
Qwen2.5-72B-Instruct 0.4982 0.3719 25.35% 0.2898 0.2733 5.68% 0.2665 0.2322 12.85% 0.3741 0.3225 13.80%
Qwen2.5-7B-Instruct 0.4524 0.3145 30.49% 0.3179 0.3090 2.81% 0.2815 0.2612 7.21% 0.2977 0.2639 11.34%
ModernBERT, Round-trip Translation, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2346 0.1468 37.42% 0.2223 0.1773 20.26% 0.2463 0.2101 14.69% 0.2178 0.1841 15.47%
gemma-2-2b-it 0.2677 0.1798 32.85% 0.2190 0.1626 25.74% 0.2474 0.2075 16.13% 0.2622 0.2088 20.39%
gemma-2-9b-it 0.2319 0.1455 37.25% 0.2156 0.1696 21.36% 0.2714 0.2255 16.93% 0.2552 0.2037 20.20%
Llama-3.1-70B-Instruct 0.3676 0.2230 39.34% 0.3063 0.2401 21.60% 0.2631 0.2195 16.56% 0.3380 0.2552 24.49%
Llama-3.1-8B-Instruct 0.3743 0.2317 38.10% 0.3304 0.2491 24.59% 0.2720 0.2136 21.47% 0.3535 0.2459 30.45%
Qwen2.5-32B-Instruct 0.3013 0.1772 41.17% 0.2589 0.1915 26.01% 0.2480 0.2108 14.98% 0.3233 0.2455 24.07%
Qwen2.5-3B-Instruct 0.3244 0.1936 40.31% 0.3226 0.2174 32.61% 0.2568 0.2062 19.71% 0.3431 0.2576 24.93%
Qwen2.5-72B-Instruct 0.4344 0.2658 38.80% 0.3276 0.2301 29.76% 0.2563 0.2176 15.11% 0.4209 0.3337 20.73%
Qwen2.5-7B-Instruct 0.3830 0.2390 37.59% 0.3266 0.2633 19.38% 0.2757 0.2262 17.93% 0.3189 0.2576 19.23%
ModernBERT, Typos, F1 Score
PopQA Natural Questions MS MARCO EntityQuestions
Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta Original Rewritten Delta
gemma-2-27b-it 0.2346 0.1890 19.43% 0.2223 0.2031 8.64% 0.2463 0.2343 4.84% 0.2178 0.1980 9.09%
gemma-2-2b-it 0.2677 0.2134 20.31% 0.2190 0.1956 10.65% 0.2474 0.2353 4.89% 0.2622 0.2312 11.85%
gemma-2-9b-it 0.2319 0.1833 20.93% 0.2156 0.1956 9.30% 0.2714 0.2552 5.97% 0.2552 0.2316 9.24%
Llama-3.1-70B-Instruct 0.3676 0.2716 26.12% 0.3063 0.2570 16.10% 0.2631 0.2466 6.27% 0.3380 0.2782 17.68%
Llama-3.1-8B-Instruct 0.3743 0.2796 25.32% 0.3304 0.2796 15.38% 0.2720 0.2529 7.03% 0.3535 0.3190 9.76%
Qwen2.5-32B-Instruct 0.3013 0.2270 24.65% 0.2589 0.2282 11.86% 0.2480 0.2357 4.95% 0.3233 0.2756 14.75%
Qwen2.5-3B-Instruct 0.3244 0.2375 26.78% 0.3226 0.2702 16.25% 0.2568 0.2398 6.63% 0.3431 0.2970 13.43%
Qwen2.5-72B-Instruct 0.4344 0.3334 23.23% 0.3276 0.2752 15.99% 0.2563 0.2433 5.06% 0.4209 0.3654 13.20%
Qwen2.5-7B-Instruct 0.3830 0.2768 27.73% 0.3266 0.2767 15.29% 0.2757 0.2575 6.59% 0.3189 0.2654 16.78%
Table 14: RAG experiment results with ModernBERT as retrieval model on F1 scores.
Appendix E Full Query Expansion Experiment Results
The following tables provide results of Retrieval and Generation components across MS MARCO and PopQA.
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
MS MARCO RTT 37.10 43.74 50.56 63.76
MS MARCO Typos 37.02 43.66 50.40 63.72
MS MARCO Readability 29.22 37.34 45.60 61.00
MS MARCO Formality 28.92 36.88 45.16 60.50
MS MARCO Politeness 29.44 37.54 45.68 60.52
PopQA RTT 64.50 69.72 74.18 83.32
PopQA Typos 64.42 69.60 73.82 83.36
PopQA Readability 64.30 69.62 74.70 84.28
PopQA Formality 72.82 77.68 81.30 88.00
PopQA Politeness 66.74 71.62 76.50 85.94
Table 15: Contriever retrieval results with RAG+HyDE for original queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
MS MARCO RTT 32.80 38.74 44.46 57.70
MS MARCO Typos 36.00 42.10 49.36 62.84
MS MARCO Readability 26.14 34.88 42.74 58.52
MS MARCO Formality 25.66 33.06 41.38 58.60
MS MARCO Politeness 28.44 36.28 44.58 59.44
PopQA RTT 44.62 49.06 53.80 65.72
PopQA Typos 59.72 64.58 69.50 80.30
PopQA Readability 59.00 63.96 68.72 79.34
PopQA Formality 68.60 72.82 76.42 84.20
PopQA Politeness 65.18 70.06 74.60 83.70
Table 16: Contriever retrieval results with RAG+HyDE for rewritten queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
PopQA RTT 0.5984 0.6496 0.6980 0.7936
PopQA Typos 0.5928 0.6444 0.6962 0.7970
PopQA Readability 0.6176 0.6720 0.7170 0.8164
PopQA Formality 0.7064 0.7518 0.7938 0.8670
PopQA Politeness 0.6322 0.6780 0.7272 0.8278
MS MARCO RTT 0.2994 0.3672 0.4362 0.5546
MS MARCO Typos 0.3030 0.3698 0.4376 0.5560
MS MARCO Readability 0.3830 0.4616 0.5414 0.6608
MS MARCO Formality 0.3816 0.4594 0.5380 0.6520
MS MARCO Politeness 0.3800 0.4580 0.5314 0.6490
Table 17: ModernBERT retrieval results with RAG+HyDE for original queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
PopQA RTT 0.3750 0.4230 0.4678 0.5802
PopQA Typos 0.5388 0.5940 0.6538 0.7584
PopQA Readability 0.5468 0.5904 0.6394 0.7402
PopQA Formality 0.6520 0.6970 0.7404 0.8228
PopQA Politeness 0.6148 0.6640 0.7094 0.8064
MS MARCO RTT 0.2438 0.3020 0.3694 0.4986
MS MARCO Typos 0.2876 0.3546 0.4280 0.5438
MS MARCO Readability 0.3462 0.4326 0.5172 0.6480
MS MARCO Formality 0.3532 0.4376 0.5122 0.6402
MS MARCO Politeness 0.3724 0.4510 0.5232 0.6454
Table 18: ModernBERT retrieval results with RAG+HyDE for rewritten queries
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.1466 0.1014 30.83%
MS MARCO RTT Qwen2.5-7B-Instruct 0.1418 0.1050 25.95%
MS MARCO RTT gemma-2-9b-it 0.1366 0.0966 29.28%
MS MARCO Typos Llama-3.1-8B-Instruct 0.1466 0.1378 6.00%
MS MARCO Typos Qwen2.5-7B-Instruct 0.1418 0.1328 6.35%
MS MARCO Typos gemma-2-9b-it 0.1366 0.1284 6.00%
MS MARCO Formality Llama-3.1-8B-Instruct 0.1626 0.1430 12.05%
MS MARCO Formality Qwen2.5-7B-Instruct 0.1556 0.1312 15.68%
MS MARCO Formality gemma-2-9b-it 0.1468 0.1296 11.72%
MS MARCO politeness Llama-3.1-8B-Instruct 0.1866 0.1782 4.50%
MS MARCO politeness Qwen2.5-7B-Instruct 0.1880 0.1750 6.91%
MS MARCO politeness gemma-2-9b-it 0.1742 0.1634 6.20%
MS MARCO Readability Llama-3.1-8B-Instruct 0.1926 0.1698 11.84%
MS MARCO Readability Qwen2.5-7B-Instruct 0.1884 0.1786 5.20%
MS MARCO Readability gemma-2-9b-it 0.1796 0.1648 8.24%
PopQA RTT Llama-3.1-8B-Instruct 0.5194 0.3242 37.58%
PopQA RTT Qwen2.5-7B-Instruct 0.5160 0.3260 36.82%
PopQA RTT gemma-2-9b-it 0.5266 0.3340 36.57%
PopQA Typos Llama-3.1-8B-Instruct 0.5194 0.4698 9.55%
PopQA Typos Qwen2.5-7B-Instruct 0.5160 0.4710 8.72%
PopQA Typos gemma-2-9b-it 0.5266 0.4842 8.05%
PopQA Formality Llama-3.1-8B-Instruct 0.6400 0.5984 6.50%
PopQA Formality Qwen2.5-7B-Instruct 0.6300 0.5908 6.22%
PopQA Formality gemma-2-9b-it 0.6348 0.5954 6.21%
PopQA politeness Llama-3.1-8B-Instruct 0.5564 0.5384 3.24%
PopQA politeness Qwen2.5-7B-Instruct 0.5364 0.5214 2.80%
PopQA politeness gemma-2-9b-it 0.5464 0.5198 4.87%
PopQA Readability Llama-3.1-8B-Instruct 0.5558 0.5086 8.49%
PopQA Readability Qwen2.5-7B-Instruct 0.5492 0.4840 11.87%
PopQA Readability gemma-2-9b-it 0.5524 0.4970 10.03%
Table 19: RAG experiment results with query expansion and Contriever as retrieval model on AM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.0318 0.0166 47.80%
MS MARCO RTT Qwen2.5-7B-Instruct 0.0358 0.0244 31.84%
MS MARCO RTT gemma-2-9b-it 0.0350 0.0254 27.43%
MS MARCO Typos Llama-3.1-8B-Instruct 0.0318 0.0286 10.06%
MS MARCO Typos Qwen2.5-7B-Instruct 0.0358 0.0294 17.88%
MS MARCO Typos gemma-2-9b-it 0.0350 0.0314 10.29%
MS MARCO Formality Llama-3.1-8B-Instruct 0.0234 0.0098 58.12%
MS MARCO Formality Qwen2.5-7B-Instruct 0.0356 0.0240 32.58%
MS MARCO Formality gemma-2-9b-it 0.0324 0.0152 53.09%
MS MARCO politeness Llama-3.1-8B-Instruct 0.0286 0.0104 63.64%
MS MARCO politeness Qwen2.5-7B-Instruct 0.0434 0.0328 24.42%
MS MARCO politeness gemma-2-9b-it 0.0462 0.0132 71.43%
MS MARCO Readability Llama-3.1-8B-Instruct 0.0182 0.0094 48.35%
MS MARCO Readability Qwen2.5-7B-Instruct 0.0374 0.0558 49.20%
MS MARCO Readability gemma-2-9b-it 0.0428 0.0554 29.44%
PopQA RTT Llama-3.1-8B-Instruct 0.2926 0.1594 45.52%
PopQA RTT Qwen2.5-7B-Instruct 0.2886 0.1628 43.59%
PopQA RTT gemma-2-9b-it 0.0076 0.0066 13.16%
PopQA Typos Llama-3.1-8B-Instruct 0.2926 0.2406 17.77%
PopQA Typos Qwen2.5-7B-Instruct 0.2886 0.2368 17.95%
PopQA Typos gemma-2-9b-it 0.0076 0.0068 10.53%
PopQA Formality Llama-3.1-8B-Instruct 0.3792 0.2488 34.39%
PopQA Formality Qwen2.5-7B-Instruct 0.4646 0.3114 32.97%
PopQA Formality gemma-2-9b-it 0.0202 0.0124 38.61%
PopQA politeness Llama-3.1-8B-Instruct 0.2424 0.1970 18.73%
PopQA politeness Qwen2.5-7B-Instruct 0.3188 0.3166 0.69%
PopQA politeness gemma-2-9b-it 0.0170 0.0058 65.88%
PopQA Readability Llama-3.1-8B-Instruct 0.2868 0.1770 38.28%
PopQA Readability Qwen2.5-7B-Instruct 0.3780 0.2338 38.15%
PopQA Readability gemma-2-9b-it 0.0114 0.0590 417.54%
Table 20: RAG experiment results with query expansion and Contriever as retrieval model on EM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.2412 0.1894 21.49%
MS MARCO RTT Qwen2.5-7B-Instruct 0.2453 0.2037 16.93%
MS MARCO RTT gemma-2-9b-it 0.2458 0.1995 18.81%
MS MARCO Typos Llama-3.1-8B-Instruct 0.2412 0.2323 3.71%
MS MARCO Typos Qwen2.5-7B-Instruct 0.2453 0.2357 3.89%
MS MARCO Typos gemma-2-9b-it 0.2458 0.2381 3.12%
MS MARCO Formality Llama-3.1-8B-Instruct 0.2593 0.2270 12.44%
MS MARCO Formality Qwen2.5-7B-Instruct 0.2680 0.2354 12.14%
MS MARCO Formality gemma-2-9b-it 0.2685 0.2353 12.39%
MS MARCO politeness Llama-3.1-8B-Instruct 0.2591 0.2323 10.34%
MS MARCO politeness Qwen2.5-7B-Instruct 0.2741 0.2516 8.21%
MS MARCO politeness gemma-2-9b-it 0.2790 0.2426 13.05%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2401 0.1955 18.56%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2618 0.2476 5.42%
MS MARCO Readability gemma-2-9b-it 0.2647 0.2431 8.19%
PopQA RTT Llama-3.1-8B-Instruct 0.3983 0.2358 40.79%
PopQA RTT Qwen2.5-7B-Instruct 0.4005 0.2398 40.12%
PopQA RTT gemma-2-9b-it 0.2372 0.1391 41.37%
PopQA Typos Llama-3.1-8B-Instruct 0.3983 0.3483 12.55%
PopQA Typos Qwen2.5-7B-Instruct 0.4005 0.3508 12.41%
PopQA Typos gemma-2-9b-it 0.2372 0.2174 8.32%
PopQA Formality Llama-3.1-8B-Instruct 0.5024 0.3968 21.02%
PopQA Formality Qwen2.5-7B-Instruct 0.5602 0.4461 20.36%
PopQA Formality gemma-2-9b-it 0.2958 0.2812 4.94%
PopQA politeness Llama-3.1-8B-Instruct 0.3779 0.3398 10.07%
PopQA politeness Qwen2.5-7B-Instruct 0.4316 0.4259 1.34%
PopQA politeness gemma-2-9b-it 0.2605 0.2277 12.59%
PopQA Readability Llama-3.1-8B-Instruct 0.4127 0.2956 28.37%
PopQA Readability Qwen2.5-7B-Instruct 0.4781 0.3552 25.70%
PopQA Readability gemma-2-9b-it 0.2633 0.2344 10.98%
Table 21: RAG experiment results with query expansion and Contriever as retrieval model on F1 scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.1706 0.1222 28.37%
MS MARCO RTT Qwen2.5-7B-Instruct 0.1634 0.1216 25.58%
MS MARCO RTT gemma-2-9b-it 0.1560 0.1138 27.05%
MS MARCO Typos Llama-3.1-8B-Instruct 0.1746 0.1642 5.96%
MS MARCO Typos Qwen2.5-7B-Instruct 0.1592 0.1522 4.40%
MS MARCO Typos gemma-2-9b-it 0.1578 0.1462 7.35%
MS MARCO Formality Llama-3.1-8B-Instruct 0.1896 0.1710 9.81%
MS MARCO Formality Qwen2.5-7B-Instruct 0.1778 0.1568 11.81%
MS MARCO Formality gemma-2-9b-it 0.1620 0.1448 10.62%
MS MARCO politeness Llama-3.1-8B-Instruct 0.2174 0.2068 4.88%
MS MARCO politeness Qwen2.5-7B-Instruct 0.2042 0.1956 4.21%
MS MARCO politeness gemma-2-9b-it 0.1910 0.1828 4.29%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2126 0.1896 10.82%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2096 0.1908 8.97%
MS MARCO Readability gemma-2-9b-it 0.1924 0.1792 6.86%
PopQA RTT Llama-3.1-8B-Instruct 0.4892 0.2890 40.92%
PopQA RTT Qwen2.5-7B-Instruct 0.4858 0.2982 38.62%
PopQA RTT gemma-2-9b-it 0.4982 0.3002 39.74%
PopQA Typos Llama-3.1-8B-Instruct 0.4910 0.4394 10.51%
PopQA Typos Qwen2.5-7B-Instruct 0.4838 0.4356 9.96%
PopQA Typos gemma-2-9b-it 0.4984 0.4524 9.23%
PopQA Formality Llama-3.1-8B-Instruct 0.6202 0.5764 7.06%
PopQA Formality Qwen2.5-7B-Instruct 0.6154 0.5718 7.08%
PopQA Formality gemma-2-9b-it 0.6168 0.5720 7.26%
PopQA politeness Llama-3.1-8B-Instruct 0.5418 0.5156 4.84%
PopQA politeness Qwen2.5-7B-Instruct 0.5242 0.5008 4.46%
PopQA politeness gemma-2-9b-it 0.5314 0.5026 5.42%
PopQA Readability Llama-3.1-8B-Instruct 0.5366 0.4662 13.12%
PopQA Readability Qwen2.5-7B-Instruct 0.5226 0.4448 14.89%
PopQA Readability gemma-2-9b-it 0.5306 0.4578 13.72%
Table 22: RAG experiment results with query expansion and ModernBERT as retrieval model on AM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.0292 0.0218 25.34%
MS MARCO RTT Qwen2.5-7B-Instruct 0.0380 0.0348 8.42%
MS MARCO RTT gemma-2-9b-it 0.0392 0.0270 31.12%
MS MARCO Typos Llama-3.1-8B-Instruct 0.0358 0.0316 11.73%
MS MARCO Typos Qwen2.5-7B-Instruct 0.0416 0.0322 22.60%
MS MARCO Typos gemma-2-9b-it 0.0400 0.0334 16.50%
MS MARCO Formality Llama-3.1-8B-Instruct 0.0288 0.0142 50.69%
MS MARCO Formality Qwen2.5-7B-Instruct 0.0396 0.0322 18.69%
MS MARCO Formality gemma-2-9b-it 0.0362 0.0154 57.46%
MS MARCO politeness Llama-3.1-8B-Instruct 0.0278 0.0116 58.27%
MS MARCO politeness Qwen2.5-7B-Instruct 0.0454 0.0400 11.89%
MS MARCO politeness gemma-2-9b-it 0.0504 0.0144 71.43%
MS MARCO Readability Llama-3.1-8B-Instruct 0.0210 0.0092 56.19%
MS MARCO Readability Qwen2.5-7B-Instruct 0.0422 0.0608 44.08%
MS MARCO Readability gemma-2-9b-it 0.0486 0.0598 23.05%
PopQA RTT Llama-3.1-8B-Instruct 0.2826 0.1408 50.18%
PopQA RTT Qwen2.5-7B-Instruct 0.2866 0.1430 50.10%
PopQA RTT gemma-2-9b-it 0.0082 0.0066 19.51%
PopQA Typos Llama-3.1-8B-Instruct 0.2768 0.2328 15.90%
PopQA Typos Qwen2.5-7B-Instruct 0.2758 0.2194 20.45%
PopQA Typos gemma-2-9b-it 0.0066 0.0086 30.30%
PopQA Formality Llama-3.1-8B-Instruct 0.3728 0.2400 35.62%
PopQA Formality Qwen2.5-7B-Instruct 0.4604 0.3054 33.67%
PopQA Formality gemma-2-9b-it 0.0162 0.0118 27.16%
PopQA politeness Llama-3.1-8B-Instruct 0.2424 0.1956 19.31%
PopQA politeness Qwen2.5-7B-Instruct 0.3240 0.3130 3.40%
PopQA politeness gemma-2-9b-it 0.0166 0.0062 62.65%
PopQA Readability Llama-3.1-8B-Instruct 0.2726 0.1550 43.14%
PopQA Readability Qwen2.5-7B-Instruct 0.3552 0.2126 40.15%
PopQA Readability gemma-2-9b-it 0.0122 0.0544 345.90%
Table 23: RAG experiment results with query expansion and ModernBERT as retrieval model on EM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.2593 0.2080 19.76%
MS MARCO RTT Qwen2.5-7B-Instruct 0.2697 0.2247 16.69%
MS MARCO RTT gemma-2-9b-it 0.2669 0.2176 18.48%
MS MARCO Typos Llama-3.1-8B-Instruct 0.2667 0.2544 4.62%
MS MARCO Typos Qwen2.5-7B-Instruct 0.2717 0.2583 4.95%
MS MARCO Typos gemma-2-9b-it 0.2677 0.2540 5.12%
MS MARCO Formality Llama-3.1-8B-Instruct 0.2855 0.2559 10.36%
MS MARCO Formality Qwen2.5-7B-Instruct 0.2939 0.2657 9.59%
MS MARCO Formality gemma-2-9b-it 0.2877 0.2528 12.14%
MS MARCO politeness Llama-3.1-8B-Instruct 0.2818 0.2558 9.22%
MS MARCO politeness Qwen2.5-7B-Instruct 0.2967 0.2745 7.49%
MS MARCO politeness gemma-2-9b-it 0.2966 0.2608 12.07%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2613 0.2135 18.29%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2825 0.2686 4.95%
MS MARCO Readability gemma-2-9b-it 0.2847 0.2634 7.47%
PopQA RTT Llama-3.1-8B-Instruct 0.3849 0.2103 45.37%
PopQA RTT Qwen2.5-7B-Instruct 0.3903 0.2185 44.01%
PopQA RTT gemma-2-9b-it 0.2275 0.1272 44.06%
PopQA Typos Llama-3.1-8B-Instruct 0.3779 0.3289 12.98%
PopQA Typos Qwen2.5-7B-Instruct 0.3798 0.3266 14.02%
PopQA Typos gemma-2-9b-it 0.2255 0.2075 8.02%
PopQA Formality Llama-3.1-8B-Instruct 0.4943 0.3861 21.89%
PopQA Formality Qwen2.5-7B-Instruct 0.5529 0.4344 21.45%
PopQA Formality gemma-2-9b-it 0.2870 0.2733 4.77%
PopQA politeness Llama-3.1-8B-Instruct 0.3719 0.3324 10.61%
PopQA politeness Qwen2.5-7B-Instruct 0.4301 0.4136 3.83%
PopQA politeness gemma-2-9b-it 0.2553 0.2230 12.65%
PopQA Readability Llama-3.1-8B-Instruct 0.3974 0.2690 32.29%
PopQA Readability Qwen2.5-7B-Instruct 0.4523 0.3272 27.66%
PopQA Readability gemma-2-9b-it 0.2557 0.2179 14.79%
Table 24: RAG experiment results with query expansion and ModernBERT as retrieval model on F1 scores.
Appendix F Full Reranking Experiment Results
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
MS MARCO RTT 0.4620 0.5208 0.5618 0.6102
MS MARCO Typos 0.4620 0.5208 0.5618 0.6102
MS MARCO Readability 0.4058 0.4688 0.5188 0.5718
MS MARCO Formality 0.4070 0.4714 0.5154 0.5598
MS MARCO Politeness 0.4078 0.4706 0.5146 0.5664
PopQA RTT 0.6640 0.7230 0.7674 0.8192
PopQA Typos 0.6640 0.7230 0.7674 0.8192
PopQA Readability 0.6864 0.7506 0.7870 0.8344
PopQA Formality 0.7620 0.8100 0.8404 0.8760
PopQA Politeness 0.7050 0.7722 0.8096 0.8534
Table 25: Contriever retrieval results with RAG+Reranking for original queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
MS MARCO RTT 0.4144 0.4674 0.5158 0.5754
MS MARCO Typos 0.4106 0.4598 0.4932 0.5440
MS MARCO Readability 0.3061 0.3773 0.4420 0.5235
MS MARCO Formality 0.2755 0.3341 0.3803 0.4365
MS MARCO Politeness 0.3665 0.4304 0.4751 0.5448
PopQA RTT 0.4900 0.5510 0.5946 0.6654
PopQA Typos 0.5420 0.5888 0.6238 0.6724
PopQA Readability 0.5839 0.6421 0.6835 0.7332
PopQA Formality 0.6837 0.7266 0.7543 0.7856
PopQA Politeness 0.6655 0.7300 0.7701 0.8081
Table 26: Contriever retrieval results with RAG+Reranking for rewritten queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
PopQA RTT 0.6836 0.7462 0.7934 0.8344
PopQA Typos 0.6836 0.7462 0.7934 0.8344
PopQA Readability 0.7012 0.7620 0.8024 0.8432
PopQA Formality 0.7684 0.8162 0.8502 0.8832
PopQA Politeness 0.7282 0.7878 0.8310 0.8688
MS MARCO RTT 0.3628 0.4302 0.4860 0.5680
MS MARCO Typos 0.3628 0.4302 0.4860 0.5680
MS MARCO Readability 0.4428 0.5194 0.5876 0.6746
MS MARCO Formality 0.4482 0.5284 0.5932 0.6720
MS MARCO Politeness 0.4478 0.5222 0.5802 0.6638
Table 27: ModernBERT retrieval results with RAG+Reranking for original queries
Dataset Linguistic Dimension R@5 R@10 R@20 R@100
PopQA RTT 0.5096 0.5744 0.6204 0.6856
PopQA Typos 0.5940 0.6496 0.6886 0.7406
PopQA Readability 0.6237 0.6836 0.7289 0.7796
PopQA Formality 0.7337 0.7809 0.8113 0.8493
PopQA Politeness 0.6888 0.7487 0.7860 0.8317
MS MARCO RTT 0.2940 0.3558 0.4160 0.5146
MS MARCO Typos 0.3248 0.3902 0.4548 0.5396
MS MARCO Readability 0.3279 0.4179 0.4994 0.6461
MS MARCO Formality 0.3479 0.4392 0.5224 0.6479
MS MARCO Politeness 0.4060 0.4882 0.5553 0.6579
Table 28: ModernBERT retrieval results with RAG+Reranking for rewritten queries
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.1428 0.1302 8.82%
MS MARCO RTT Qwen2.5-7B-Instruct 0.1416 0.1290 8.90%
MS MARCO RTT gemma-2-9b-it 0.1292 0.1202 6.97%
MS MARCO Typos Llama-3.1-8B-Instruct 0.1428 0.1118 21.71%
MS MARCO Typos Qwen2.5-7B-Instruct 0.1416 0.1120 20.90%
MS MARCO Typos gemma-2-9b-it 0.1292 0.1066 17.49%
MS MARCO Formality Llama-3.1-8B-Instruct 0.1552 0.1127 27.36%
MS MARCO Formality Qwen2.5-7B-Instruct 0.1540 0.1100 28.57%
MS MARCO Formality gemma-2-9b-it 0.1430 0.0979 31.56%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.1876 0.1586 15.46%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.1820 0.1573 13.55%
MS MARCO Politeness gemma-2-9b-it 0.1730 0.1489 13.95%
MS MARCO Readability Llama-3.1-8B-Instruct 0.1820 0.1528 16.04%
MS MARCO Readability Qwen2.5-7B-Instruct 0.1780 0.1517 14.76%
MS MARCO Readability gemma-2-9b-it 0.1630 0.1449 11.12%
PopQA RTT Llama-3.1-8B-Instruct 0.5184 0.3366 35.07%
PopQA RTT Qwen2.5-7B-Instruct 0.4924 0.3350 31.97%
PopQA RTT gemma-2-9b-it 0.5086 0.3504 31.10%
PopQA Typos Llama-3.1-8B-Instruct 0.5184 0.3490 32.68%
PopQA Typos Qwen2.5-7B-Instruct 0.4924 0.3352 31.93%
PopQA Typos gemma-2-9b-it 0.5086 0.3660 28.04%
PopQA Formality Llama-3.1-8B-Instruct 0.6554 0.5099 22.20%
PopQA Formality Qwen2.5-7B-Instruct 0.6260 0.4861 22.35%
PopQA Formality gemma-2-9b-it 0.6294 0.4999 20.58%
PopQA Politeness Llama-3.1-8B-Instruct 0.5144 0.4811 6.48%
PopQA Politeness Qwen2.5-7B-Instruct 0.4818 0.4540 5.77%
PopQA Politeness gemma-2-9b-it 0.4978 0.4685 5.88%
PopQA Readability Llama-3.1-8B-Instruct 0.5142 0.4158 19.14%
PopQA Readability Qwen2.5-7B-Instruct 0.4870 0.3874 20.45%
PopQA Readability gemma-2-9b-it 0.5028 0.4091 18.63%
Table 29: Generation results with Contriever retrieval and reranking on AM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.0288 0.0214 25.69%
MS MARCO RTT Qwen2.5-7B-Instruct 0.0336 0.0302 10.12%
MS MARCO RTT gemma-2-9b-it 0.0388 0.0288 25.77%
MS MARCO Typos Llama-3.1-8B-Instruct 0.0288 0.0174 39.58%
MS MARCO Typos Qwen2.5-7B-Instruct 0.0336 0.0254 24.40%
MS MARCO Typos gemma-2-9b-it 0.0388 0.0308 20.62%
MS MARCO Formality Llama-3.1-8B-Instruct 0.0232 0.0059 74.43%
MS MARCO Formality Qwen2.5-7B-Instruct 0.0362 0.0149 58.93%
MS MARCO Formality gemma-2-9b-it 0.0344 0.0090 73.84%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.0248 0.0079 68.28%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.0384 0.0265 30.90%
MS MARCO Politeness gemma-2-9b-it 0.0512 0.0153 70.18%
MS MARCO Readability Llama-3.1-8B-Instruct 0.0166 0.0076 54.22%
MS MARCO Readability Qwen2.5-7B-Instruct 0.0330 0.0441 33.74%
MS MARCO Readability gemma-2-9b-it 0.0452 0.0521 15.34%
PopQA RTT Llama-3.1-8B-Instruct 0.2616 0.1454 44.42%
PopQA RTT Qwen2.5-7B-Instruct 0.2646 0.1646 37.79%
PopQA RTT gemma-2-9b-it 0.0092 0.0090 2.17%
PopQA Typos Llama-3.1-8B-Instruct 0.2616 0.1660 36.54%
PopQA Typos Qwen2.5-7B-Instruct 0.2646 0.1536 41.95%
PopQA Typos gemma-2-9b-it 0.0092 0.0094 2.17%
PopQA Formality Llama-3.1-8B-Instruct 0.3452 0.1375 60.18%
PopQA Formality Qwen2.5-7B-Instruct 0.4498 0.1851 58.86%
PopQA Formality gemma-2-9b-it 0.0218 0.0068 68.81%
PopQA Politeness Llama-3.1-8B-Instruct 0.2016 0.1649 18.22%
PopQA Politeness Qwen2.5-7B-Instruct 0.2734 0.2725 0.32%
PopQA Politeness gemma-2-9b-it 0.0166 0.0063 62.25%
PopQA Readability Llama-3.1-8B-Instruct 0.2460 0.1547 37.13%
PopQA Readability Qwen2.5-7B-Instruct 0.3212 0.1873 41.70%
PopQA Readability gemma-2-9b-it 0.0134 0.0383 185.57%
Table 30: Generation results with Contriever retrieval and reranking on EM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.2310 0.2146 7.12%
MS MARCO RTT Qwen2.5-7B-Instruct 0.2402 0.2272 5.42%
MS MARCO RTT gemma-2-9b-it 0.2394 0.2211 7.64%
MS MARCO Typos Llama-3.1-8B-Instruct 0.2310 0.2015 12.76%
MS MARCO Typos Qwen2.5-7B-Instruct 0.2402 0.2090 12.99%
MS MARCO Typos gemma-2-9b-it 0.2394 0.2124 11.26%
MS MARCO Formality Llama-3.1-8B-Instruct 0.2493 0.1913 23.26%
MS MARCO Formality Qwen2.5-7B-Instruct 0.2604 0.1984 23.81%
MS MARCO Formality gemma-2-9b-it 0.2621 0.1990 24.06%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.2494 0.2142 14.11%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.2605 0.2303 11.58%
MS MARCO Politeness gemma-2-9b-it 0.2729 0.2272 16.74%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2303 0.1816 21.15%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2491 0.2216 11.05%
MS MARCO Readability gemma-2-9b-it 0.2564 0.2255 12.04%
PopQA RTT Llama-3.1-8B-Instruct 0.3757 0.2295 38.90%
PopQA RTT Qwen2.5-7B-Instruct 0.3730 0.2440 34.57%
PopQA RTT gemma-2-9b-it 0.2291 0.1488 35.02%
PopQA Typos Llama-3.1-8B-Instruct 0.3757 0.2501 33.43%
PopQA Typos Qwen2.5-7B-Instruct 0.3730 0.2395 35.80%
PopQA Typos gemma-2-9b-it 0.2291 0.1606 29.88%
PopQA Formality Llama-3.1-8B-Instruct 0.4826 0.2859 40.75%
PopQA Formality Qwen2.5-7B-Instruct 0.5492 0.3202 41.69%
PopQA Formality gemma-2-9b-it 0.2936 0.2278 22.42%
PopQA Politeness Llama-3.1-8B-Instruct 0.3298 0.2906 11.90%
PopQA Politeness Qwen2.5-7B-Instruct 0.3804 0.3633 4.50%
PopQA Politeness gemma-2-9b-it 0.2345 0.2032 13.38%
PopQA Readability Llama-3.1-8B-Instruct 0.3691 0.2563 30.56%
PopQA Readability Qwen2.5-7B-Instruct 0.4162 0.2883 30.74%
PopQA Readability gemma-2-9b-it 0.2392 0.1888 21.06%
Table 31: Generation results with Contriever retrieval and reranking on F1 scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.1964 0.1390 29.23%
MS MARCO RTT Qwen2.5-7B-Instruct 0.1882 0.1354 28.06%
MS MARCO RTT gemma-2-9b-it 0.1690 0.1246 26.27%
MS MARCO Typos Llama-3.1-8B-Instruct 0.1964 0.1772 9.78%
MS MARCO Typos Qwen2.5-7B-Instruct 0.1882 0.1678 10.84%
MS MARCO Typos gemma-2-9b-it 0.1690 0.1618 4.26%
MS MARCO Formality Llama-3.1-8B-Instruct 0.2118 0.1655 21.84%
MS MARCO Formality Qwen2.5-7B-Instruct 0.2012 0.1531 23.89%
MS MARCO Formality gemma-2-9b-it 0.1770 0.1388 21.58%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.2394 0.2134 10.86%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.2316 0.2067 10.77%
MS MARCO Politeness gemma-2-9b-it 0.2160 0.1943 10.06%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2324 0.1791 22.92%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2254 0.1853 17.78%
MS MARCO Readability gemma-2-9b-it 0.2082 0.1702 18.25%
PopQA RTT Llama-3.1-8B-Instruct 0.5340 0.3540 33.71%
PopQA RTT Qwen2.5-7B-Instruct 0.5130 0.3476 32.24%
PopQA RTT gemma-2-9b-it 0.5264 0.3640 30.85%
PopQA Typos Llama-3.1-8B-Instruct 0.5340 0.4546 14.87%
PopQA Typos Qwen2.5-7B-Instruct 0.5130 0.4348 15.24%
PopQA Typos gemma-2-9b-it 0.5264 0.4648 11.70%
PopQA Formality Llama-3.1-8B-Instruct 0.6572 0.6163 6.22%
PopQA Formality Qwen2.5-7B-Instruct 0.6238 0.5969 4.32%
PopQA Formality gemma-2-9b-it 0.6336 0.5979 5.63%
PopQA Politeness Llama-3.1-8B-Instruct 0.5686 0.5427 4.55%
PopQA Politeness Qwen2.5-7B-Instruct 0.5400 0.5093 5.68%
PopQA Politeness gemma-2-9b-it 0.5484 0.5187 5.41%
PopQA Readability Llama-3.1-8B-Instruct 0.5694 0.4929 13.44%
PopQA Readability Qwen2.5-7B-Instruct 0.5390 0.4661 13.53%
PopQA Readability gemma-2-9b-it 0.5532 0.4800 13.23%
Table 32: Generation results with ModernBERT retrieval and reranking on AM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.0446 0.0248 44.39%
MS MARCO RTT Qwen2.5-7B-Instruct 0.0472 0.0370 21.61%
MS MARCO RTT gemma-2-9b-it 0.0428 0.0326 23.83%
MS MARCO Typos Llama-3.1-8B-Instruct 0.0446 0.0342 23.32%
MS MARCO Typos Qwen2.5-7B-Instruct 0.0472 0.0376 20.34%
MS MARCO Typos gemma-2-9b-it 0.0428 0.0352 17.76%
MS MARCO Formality Llama-3.1-8B-Instruct 0.0332 0.0105 68.27%
MS MARCO Formality Qwen2.5-7B-Instruct 0.0496 0.0240 51.61%
MS MARCO Formality gemma-2-9b-it 0.0366 0.0087 76.14%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.0348 0.0131 62.45%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.0572 0.0371 35.20%
MS MARCO Politeness gemma-2-9b-it 0.0550 0.0141 74.30%
MS MARCO Readability Llama-3.1-8B-Instruct 0.0214 0.0091 57.63%
MS MARCO Readability Qwen2.5-7B-Instruct 0.0456 0.0566 24.12%
MS MARCO Readability gemma-2-9b-it 0.0492 0.0559 13.69%
PopQA RTT Llama-3.1-8B-Instruct 0.2626 0.1582 39.76%
PopQA RTT Qwen2.5-7B-Instruct 0.2794 0.1710 38.80%
PopQA RTT gemma-2-9b-it 0.0120 0.0102 15.00%
PopQA Typos Llama-3.1-8B-Instruct 0.2626 0.2084 20.64%
PopQA Typos Qwen2.5-7B-Instruct 0.2794 0.2166 22.48%
PopQA Typos gemma-2-9b-it 0.0120 0.0122 1.67%
PopQA Formality Llama-3.1-8B-Instruct 0.3422 0.1481 56.71%
PopQA Formality Qwen2.5-7B-Instruct 0.4472 0.2155 51.82%
PopQA Formality gemma-2-9b-it 0.0244 0.0073 69.95%
PopQA Politeness Llama-3.1-8B-Instruct 0.2006 0.1761 12.23%
PopQA Politeness Qwen2.5-7B-Instruct 0.3020 0.3095 2.49%
PopQA Politeness gemma-2-9b-it 0.0216 0.0083 61.73%
PopQA Readability Llama-3.1-8B-Instruct 0.2526 0.1645 34.89%
PopQA Readability Qwen2.5-7B-Instruct 0.3580 0.2152 39.89%
PopQA Readability gemma-2-9b-it 0.0182 0.0479 163.37%
Table 33: Generation results with ModernBERT retrieval and reranking on EM scores.
Dataset Linguistics LLMs Original Rewritten Delta
MS MARCO RTT Llama-3.1-8B-Instruct 0.2837 0.2203 22.35%
MS MARCO RTT Qwen2.5-7B-Instruct 0.2866 0.2349 18.04%
MS MARCO RTT gemma-2-9b-it 0.2797 0.2308 17.49%
MS MARCO Typos Llama-3.1-8B-Instruct 0.2837 0.2622 7.56%
MS MARCO Typos Qwen2.5-7B-Instruct 0.2866 0.2672 6.78%
MS MARCO Typos gemma-2-9b-it 0.2797 0.2638 5.68%
MS MARCO Formality Llama-3.1-8B-Instruct 0.3004 0.2395 20.25%
MS MARCO Formality Qwen2.5-7B-Instruct 0.3094 0.2477 19.96%
MS MARCO Formality gemma-2-9b-it 0.3011 0.2440 18.95%
MS MARCO Politeness Llama-3.1-8B-Instruct 0.3011 0.2602 13.58%
MS MARCO Politeness Qwen2.5-7B-Instruct 0.3139 0.2778 11.51%
MS MARCO Politeness gemma-2-9b-it 0.3173 0.2664 16.02%
MS MARCO Readability Llama-3.1-8B-Instruct 0.2736 0.2099 23.27%
MS MARCO Readability Qwen2.5-7B-Instruct 0.2925 0.2595 11.28%
MS MARCO Readability gemma-2-9b-it 0.2935 0.2581 12.04%
PopQA RTT Llama-3.1-8B-Instruct 0.3850 0.2438 36.67%
PopQA RTT Qwen2.5-7B-Instruct 0.3903 0.2531 35.15%
PopQA RTT gemma-2-9b-it 0.2396 0.1546 35.49%
PopQA Typos Llama-3.1-8B-Instruct 0.3850 0.3173 17.57%
PopQA Typos Qwen2.5-7B-Instruct 0.3903 0.3175 18.65%
PopQA Typos gemma-2-9b-it 0.2396 0.2090 12.76%
PopQA Formality Llama-3.1-8B-Instruct 0.4808 0.3310 31.16%
PopQA Formality Qwen2.5-7B-Instruct 0.5468 0.3868 29.27%
PopQA Formality gemma-2-9b-it 0.2973 0.2773 6.73%
PopQA Politeness Llama-3.1-8B-Instruct 0.3466 0.3192 7.91%
PopQA Politeness Qwen2.5-7B-Instruct 0.4185 0.4083 2.44%
PopQA Politeness gemma-2-9b-it 0.2617 0.2258 13.69%
PopQA Readability Llama-3.1-8B-Instruct 0.3906 0.2886 26.12%
PopQA Readability Qwen2.5-7B-Instruct 0.4571 0.3353 26.64%
PopQA Readability gemma-2-9b-it 0.2671 0.2240 16.13%
Table 34: Generation results with ModernBERT retrieval and reranking on F1 scores.
Appendix G Rewriting Prompts
G.1 Formality
Prompt 1:
You are an AI assistant skilled at transforming formal queries into casual, everyday language. Rewrite the following query so that it sounds very informal. Experiment with different colloquial openings, varied sentence constructions, and a mix of slang, idioms, and casual expressions throughout the sentence. Avoid using the same phrase repeatedly (e.g., "hey, so like") and ensure the meaning remains unchanged.
Prompt 2:
Your task is to convert the given query into an informal version that feels natural and conversational. Instead of a uniform introductory phrase, use a range of informal expressions (such as interjections, casual questions, or slang) at different parts of the sentence. Mix up the structure—sometimes start with an interjection, other times rephrase the sentence completely—while keeping the original meaning intact.
Prompt 3:
Task: Transform Formal to Extremely Informal Language
Convert the following formal sentence into an extremely informal, messy, and natural version. The output should sound like authentic, real-world casual speech—as if spoken in an informal chat, online conversation, or street talk.
Critical Rules: - Diversity is key: No two sentences should follow the same pattern. - Break formal sentence structures: Chop up long phrases, reorder words, or make them flow casually. - Use a wide range of informality techniques—DO NOT rely on just contractions or slang. - Avoid starting every sentence with the same phrase or discourse marker. - Embrace randomness: Let the outputs sound wild, real, and unpredictable. - Do not start your sentences with "Yo", "Hey so like", etc.
Final Execution Instruction: Generate an informal version of the following sentence that: - Uses multiple different informality techniques. - Avoids repetitive sentence structures or patterns. - Sounds raw, conversational, and unpredictable.
Original Query:
G.2 Readablility
Prompt 1:
1. Task Definition:
You are rewriting a query to make it significantly less readable while preserving the original semantic meaning as closely as possible.
2. Constraints & Goals:
- Flesch Reading Ease Score: The rewritten text must have a Flesch score below 60 (preferably below 50).
- Semantic Similarity: The rewritten text must have SBERT similarity > 0.7 compared with the original query.
- Length: The rewritten text must remain approximately the same length as the original query (±10%).
- Preserve Domain Terminology: Do not remove or drastically change domain-specific words, abbreviations, or technical terms (e.g., "IRS," "distance," etc.).
- Abbreviation: Do not expand abbreviations unless the original query already used the expanded form.
- No New Information: You must not add additional details beyond what the original query states.
- Question Format: Retain the form of a question if the original is posed as a question.
3. How to Increase Complexity:
- Lexical Changes: Use advanced or academic synonyms only for common words. For domain or key terms (e.g., "distance," "IRS," "tax"), keep the original term or use a very close synonym if necessary to maintain meaning.
- Syntactic Complexity: Introduce passive voice, nominalizations, embedded clauses, and parenthetical or subordinate phrases. Ensure the sentence flow is more formal and convoluted without changing the core meaning.
- Redundancy & Formality: Employ circumlocution and excessively formal expressions (e.g., "due to the fact that" instead of "because") while avoiding any semantic drift.
- Dense, Indirect Construction: Favor longer phrases, indirect references, and wordiness. Avoid direct or simple phrasing.
Original Query: ...
Less Readable Query:
Prompt 2:
**Task Description:**
Transform a given query into a significantly less readable version while preserving its original semantic meaning as closely as possible.
**Constraints & Goals:**
- **Readability:** The rewritten text must have a **Flesch Reading Ease Score below 60**, preferably below 50.
- **Semantic Similarity:** The rewritten text must achieve an **SBERT similarity score > 0.7** with the original query.
- **Length Consistency:** The modified text should be **within ±10% of the original length**.
- **Preserve Key Terminology:** **Do not alter domain-specific words, abbreviations, or technical jargon** (e.g., "IRS," "distance").
- **Abbreviation Handling:** **Do not expand abbreviations** unless they are already expanded in the original query.
- **Maintain Original Intent:** Do not add, remove, or alter the factual content of the query.
- **Retain Question Structure:** If the input is a question, the output must also be a question.
**Techniques to Decrease Readability:**
1. **Lexical Complexity:** Replace common words with **advanced, academic, or formal synonyms**, while keeping domain-specific terms unchanged.
2. **Syntactic Complexity:** Introduce **passive voice, nominalizations, embedded clauses, or subordinate structures** to increase sentence density.
3. **Redundancy & Formality:** Use **circumlocution, excessive formality, and indirect phrasing** (e.g., "in light of the fact that" instead of "because").
4. **Dense Sentence Structure:** Prefer **wordy, indirect, and convoluted constructions** over direct phrasing.
Original Query: ...
Less Readable Query:
Prompt 3:
**Objective:**
You are tasked with **rewriting a given query to make it significantly less readable** while preserving its original semantic meaning with high fidelity.
**Guiding Principles:**
- **Readability Constraint:** The rewritten text must have a **Flesch Reading Ease Score of <= 60**, preferably <= 50.
- **Semantic Integrity:** Ensure an **SBERT similarity score of at least 0.7** between the original and rewritten text.
- **Length Tolerance:** Maintain an **approximate length deviation of no more than ±10%** from the original.
- **Terminology Preservation:** Domain-specific terms (e.g., "IRS," "distance") **must remain intact** or be substituted only with **near-synonymous equivalents**.
- **Abbreviation Handling:** If an abbreviation exists, **retain it as is** unless the original query explicitly expands it.
- **Strict Content Preservation:** Do **not introduce any new information** or omit existing details.
- **Question Retention:** If the input is a question, the reformulated output **must remain a question**.
**Techniques for Readability Reduction:**
- **Lexical Sophistication:** Replace commonplace words with **more complex, formal, or technical alternatives** while maintaining clarity of meaning.
- **Structural Density:** Employ **passive constructions, embedded clauses, and nominalized phrases** to increase syntactic complexity.
- **Circumlocution & Wordiness:** Favor **verbose, indirect expressions** over concise phrasing (e.g., "with regard to" instead of "about").
- **Elaborate Phrasing:** Use **multi-clause structures and intricate sentence formations** to reduce direct readability.
Original Query: ...
Less Readable Query:
G.3 Politeness
Prompt 1:
Task: Rewrite Queries to Sound More Polite and Courteous
Rephrase the given query into a more polite, respectful, and considerate version while preserving its original intent. The output should reflect a natural, well-mannered tone suitable for professional or friendly interactions. The generated query should be a single sentence.
Critical Rules:
- Use a variety of politeness techniques, including warm greetings, indirect requests, and expressions of gratitude.
- Avoid robotic or overly formal constructions—make it sound naturally courteous, warm and friendly.
- Do not always start your sentence with ’Could you please tell’. Use emotional undertones and specific attempts at politeness.
- Maintain the original meaning without unnecessary embellishment.
- Do not start the generated query with ’I hope you are ...’ or end with a single ’Thank you’ sentence. Generate only a single polite query sentence.
Original Query: ...
Polite Query:
Prompt 2:
Task: Enhance the Courtesy of a Given Query
Transform the provided query into a more respectful, friendly, and warm version, ensuring it conveys respect and warmth while keeping the original intent intact. The reworded request should sound engaging, professional, and well-mannered. The generated query should be a single sentence.
Key Considerations:
- Use a mix of politeness techniques, including indirect phrasing, friendly introductions, and appreciative language.
- Keep the tone natural—avoid overly rigid or formal wording that feels robotic.
- Vary sentence structures instead of defaulting to "Could you please...". Use emotional undertones and specific attempts at politeness.
- Maintain the original meaning while subtly enhancing the request’s politeness and friendliness.
- Avoid beginning the generated query with ’I hope you are...’ or concluding it with a separate ’Thank you.’ sentence. Generate only one polite query sentence.
Original Query: ...
Polite Query:
Prompt 3:
Task: Refining Queries for Politeness and Warmth
Transform a given query into a more courteous, engaging, and warm request while ensuring it retains the original intent. The revised version should sound friendly, professional, and respectful. The generated query should be a single sentence.
Guidelines:
- Incorporate politeness techniques such as indirect requests, warm introductions, and appreciative language.
- Ensure the tone is natural—avoid excessive formality that feels robotic.
- Diversify sentence structures rather than defaulting to "Could you please...". Use emotional undertones and specific attempts at politeness.
- Subtly enhance warmth and professionalism while preserving clarity and intent.
- Avoid beginning the generated query with ’I hope you are ...’ or concluding it with a standalone ’Thank you’ sentence. Generate only one polite query sentence.
Original Query: ...
Polite Query:
Appendix H LLMs Prompts
H.1 Few-shot Prompts
Few-shot examples:
PopQA:
- readability:
Question: What genre is Golden?
Answer: rock music
Question: In which specific genre does the work titled "Golden" find its classification?
Answer: rock music
- politeness:
Question: What genre is Golden?
Answer: rock music
Question: Would you be so kind as to share with me what genre Golden falls under?
Answer: rock music
- formality:
Question: What genre is Golden?
Answer: rock music
Question: Hey, so like, do you know what genre Golden is?
Answer: rock music
- round-trip translation:
Question: What genre is Golden?
Answer: rock music
Question: What genre of Golden?
Answer: rock music
- typos:
Question: What genre is Golden?
Answer: rock music
Question: What genra is Golden?
Answer: rock music
EntityQuestions:
- readability
Question: Where was Michael Jack born?
Answer: Folkestone
Question: In what geographical locale did the individual known as Michael Jackson enter into existence?
Answer: Folkestone
- politeness:
Question: Where was Michael Jack born?
Answer: Folkestone
Question: Would you be so kind as to share the birthplace of Michael Jack?
Answer: Folkestone
- formality:
Question: Where was Michael Jack born?
Answer: Folkestone
Question: Hey, so like, do you know where Michael Jack was born?
Answer: Folkestone
- round-trip translation:
Question: Where was Michael Jack born?
Answer: Folkestone
Question: Where was Michael Jacques born?
Answer: Folkestone
- typos:
Question: Where was Michael Jack born?
Answer: Folkestone
Question: Where was Michael Jack born?
Answer: Folkestone
MS MARCO:
- readability:
Question: how long can chicken stay good in the fridge
Answer: 1 to 2 days
Question: What is the time span within which chicken can sustain its quality for consumption when preserved in a refrigerated setting?
Answer: 1 to 2 days
- politeness:
Question: how long can chicken stay good in the fridge
Answer: 1 to 2 days
Question: Would you be so kind as to share how long chicken remains fresh in the refrigerator?
Answer: 1 to 2 days
- formality:
Question: how long can chicken stay good in the fridge
Answer: 1 to 2 days
Question: Hey, so like, do you know how long chicken can last in the fridge?
Answer: 1 to 2 days
- round-trip translation:
Question: how long can chicken stay good in the fridge
Answer: 1 to 2 days
Question: How long will chicken stay fresh in the refrigerator
Answer: 1 to 2 days
- typos:
Question: how long can chicken stay good in the fridge
Answer: 1 to 2 days
Question: how leng can chickon stay good in the fridge
Answer: 1 to 2 days
Natural Questions:
- readability:
Question: how many pieces in a terry’s chocolate orange
Answer: six
Question: What is the total quantity of individual segments contained within a Terry’s chocolate orange confectionery item?
Answer: six
- politeness:
Question: how many pieces in a terry’s chocolate orange
Answer: six
Question: Would you be so kind as to share the number of segments typically found in a Terry’s chocolate orange?
Answer: six
- formality:
Question: how many pieces in a terry’s chocolate orange
Answer: six
Question: Hey, so like, do you know a terry’s chocolate orange contains how many pieces
Answer: six
- round-trip translation:
Question: how many pieces in a terry’s chocolate orange
Answer: six
Question: How many pieces of Terry’s Chocolate Orange
Answer: six
- typos:
Question: how many pieces in a terry’s chocolate orange
Answer: six
Question: how meny pieces in a tarry’s chocolate orange
Answer: six
Prompt:
You are a professional question-answer task assistant. Use the following pieces of retrieved context to answer the question briefly.
Context:
contexts
Below are examples of questions and answers:
few_shot_examples
Now, it’s your turn to answer the question below. The answer should contain ONLY one sentence and DO NOT explain reasons.
Appendix I Rewriting Examples
Category Type Query
RTT Original What type of music does The Eruption of Mount St. Helens! play?
Rewritten What music the eruption of Mount St Helens! play?
RTT Original Who is Hilde Coppi married to?
Rewritten With whom was Hilde Coppi married?
RTT Original Which company is HMS Blankney produced by?
Rewritten What company is producing HMS Blankey?
RTT Original Where is Flemington Racecourse located?
Rewritten Where is Flemington Racecourse?
Typos Original Where was R. Kent Greenawalt born?
Rewritten Wher was R. Kent Greenawalt born?
Typos Original What kind of work does M. Ramanathan do?
Rewritten What kind ofh work does M. Ramanathan do?
Typos Original What type of music does El Cantor del circo play?
Rewritten What type of music does El Cantorh del circo pay?
Typos Original What is Rembrandt famous for?
Rewritten What is Rembrandt faumous for?
Formality Original Which company is Galaxy Camera produced by?
Rewritten Hey, quick question! Which company actually makes the Galaxy Camera?
Formality Original Who is the author of Intensity?
Rewritten Yo, do you know who wrote Intensity?
Formality Original Which country is Parchliny located in?
Rewritten Hey, just curious, do you know what country Parchliny is in?
Formality Original Who is Liu Bei’s child?
Rewritten Hey, so, do you know who Liu Bei’s kid is? I’m super curious about it!
Readability Original What type of music does Anbe Sivam play?
Rewritten What genre of musical compositions is performed by Anbe Sivam?
Readability Original Where was FC Utrecht founded?
Rewritten In what location was the establishment of FC Utrecht initiated?
Readability Original Where was John Ernle educated?
Rewritten At which institution did John Ernle receive his education?
Readability Original Where was The Shiru Group founded?
Rewritten In which geographical location did The Shiru Group originate?
Politeness Original What music label is Time in Place represented by?
Rewritten May I kindly inquire which music label represents Time in Place?
Politeness Original Which country was The Border Blasters created in?
Rewritten Would you be so kind as to share which country The Border Blasters originated from?
Politeness Original Which country is Oleksin, Otwock County located in?
Rewritten Could you kindly share which country Oleksin, Otwock County is situated in?
Politeness Original Where did Wolfe Tone die?
Rewritten Would you be so kind as to share the location where Wolfe Tone passed away?
Table 35: Rewriting examples across all linguistic variations from the EntityQuestions dataset, with queries split across rows for readability.

### v2

@@
 [^1] Code is available at https://github.com/Springcty/RAG-fragility-to-linguistic-variation.
+Out of Style: RAG’s Fragility to Linguistic Variation
+Tianyu Cao1thanks: Equal contribution.
+Neel Bhandari1∗   Akhila Yerukola1   Akari Asai12   Maarten Sap1 1 Language Technologies Institute, Carnegie Mellon University 2 Allen Institute for AI {tianyuca, neelbhan, ayerukol, aasai, msap2}@cs.cmu.edu
 1 Introduction
 Retrieval-augmented Generation (RAG) systems enhance Large Language Models (LLMs) by integrating external knowledge retrieval, grounding their output in factual context to improve accuracy and reliability (Lewis et al., 2021; Gao et al., 2024).
-However, their widespread integration into real-world applications (K2view, 2024) introduce potential challenges regarding robustness to linguistic variations.
-Users bring varied backgrounds, domains, and cultural contexts that naturally produce linguistic differences in their queries (Park et al., 2024; Li et al., 2020; Lorenzo-Dus & Bou-Franch, 2013).
+However, their widespread integration into real-world applications (K2view, 2024) introduces potential challenges regarding robustness to linguistic variations.
+Users bring varied backgrounds, domains, and cultural contexts that naturally produce linguistic differences in their queries (Park et al., 2024; Li et al., 2020; Lorenzo-Dus and Bou-Franch, 2013).
 As Figure 1 illustrates, different from the carefully curated queries from traditional NLP benchmarks, real-world user-LLM queries tend to be less formal and frequently contain grammatical inconsistencies (Ouyang et al., 2023).
@@
 We present a large-scale systematic investigation of how variations in linguistic characteristics affect the robustness of RAG systems.
-We focus on formality, readability, politeness, and grammatical correctness - fundamental aspects of linguistic variations that are critical and common in real-world user-LLM queries (Li et al., 2024; Yin et al., 2024; Hagen et al., 2024).
+We target diverse and prevalent variations commonly found in real-world user inputs that meaningfully challenge RAG systems (Park et al., 2024; Ouyang et al., 2023), namely, formality, readability, politeness, and grammatical correctness.
+These choices ensure our analysis covers stylistic, pragmatic, and structural aspects aligned with practical usage.
 By automatically rewriting queries across these dimensions, we analyze how linguistic variations impact each RAG system component, as well as potential cascading errors throughout the pipeline.
-Our evaluation encompasses two retrieval models, namely Contriever (Izacard et al., 2021) and ModernBERT (Nussbaum et al., 2024), and nine LLMs from three families (Llama 3.1, Qwen 2.5, Gemma 2) of varying scales across four open-domain Question Answering (QA) datasets: PopQA (Mallen et al., 2023), EntityQuestions (Sciavolino et al., 2022), MS MARCO (Bajaj et al., 2018), and Natural Questions (Kwiatkowski et al., 2019a).
+Our evaluation encompasses two retrieval models, namely Contriever (Izacard et al., 2021) and ModernBERT (Nussbaum et al., 2024), and nine LLMs from three families (Llama 3.1, Qwen 2.5, Gemma 2) of varying scales across four open-domain Question Answering (QA) datasets: PopQA (Mallen et al., 2023), EntityQuestions (Sciavolino et al., 2022), MS MARCO (Bajaj et al., 2018), and Natural Questions (Kwiatkowski et al., 2019).
 Our experiments reveal significant vulnerabilities in RAG systems to linguistic variations.
@@
 These findings highlight the urgent need to improve the robustness of retrieval components when handling linguistically varied queries.
-We also find that while advanced RAG techniques such as query expansion with HyDE (Gao et al., 2022) and documents reranking tend to improve overall performance, they remain similarly vulnerable to linguistic variations.
+We also find that while advanced RAG techniques, e.g., query expansion with HyDE (Gao et al., 2022) and documents reranking, tend to improve overall performance, they remain similarly vulnerable to linguistic variations.
 In summary, we present the first systematic analysis of the robustness of RAG systems to linguistic query reformulation.
@@
 Robustness of retrieval systems to linguistic variations.
-Prior research investigated retrieval system robustness to noisy queries (Campos et al., 2023; Chen et al., 2022; 2023b) and specific linguistic variations including word substitutions (Wu et al., 2022), aspect changes and paraphrasing (Penha et al., 2022), typographical errors (Zhuang & Zuccon, 2021), and grammatical variations (Long et al., 2024).
+Prior research investigated retrieval system robustness to noisy queries (Campos et al., 2023; Chen et al., 2022, 2023b) and specific linguistic variations including word substitutions (Wu et al., 2022), aspect changes and paraphrasing (Penha et al., 2022), typographical errors (Zhuang and Zuccon, 2021), and grammatical variations (Long et al., 2024).
 Our work provides the first holistic evaluation of RAG systems’ robustness to diverse linguistic variations and uncovers cascading failures across the entire RAG pipeline.
 Robustness of language models to linguistic variations.
-Prior research has examined impacts of syntactic perturbations Moradi & Samwald (2021); Singh et al.
-(2024), round-trip translation Bhandari & Chen (2023), politeness variations Yin et al.
+Prior research has examined impacts of syntactic perturbations Moradi and Samwald (2021); Singh et al.
+(2024), round-trip translation Bhandari and Chen (2023), politeness variations Yin et al.
 (2024), equivalent queries Cao et al.
@@
 (2025) analyzed how spurious features affect RAG perform.
-These studies focus primarily on retrieved content noise rather than initial query variations.
-Our work uniquely demonstrates how diverse linguistic variations in user queries compound throughout the RAG pipeline, exposing critical vulnerabilities in systems serving diverse users.
+Cho et al.
+(2024) introduced document-level perturbations and evaluated RAG’s vulnerability to noisy documents.
+RbFT (Tu et al., 2025) introduce specialized training objectives to improve consistency under corrupted or adversarial retrieval contexts.
+These studies focus primarily on retrieved content noise rather than initial query; our work uniquely demonstrates how diverse linguistic variations in user queries compound throughout the RAG pipeline, exposing critical vulnerabilities in systems serving diverse users.
 3 Robustness Evaluation Approach
@@
 Given a seed dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ and $y_{i}$ indicate $i$-th input and output, respectively, we reformulate each query $x_{i}\rightarrow x_{i}^{\prime}$ based on four linguistic aspects.
-A robust RAG system, composed of a retriever $\mathcal{R}$ and a generator $\mathcal{G}$ operating on corpus $\mathcal{C}$, should maintain performance when processing linguistically varied inputs.
+A robust RAG system, composed of a retriever $\mathcal{R}$ and a generator $\mathcal{G}$ operating on corpus $\mathcal{C}$, should maintain performance when processing linguistically varied inputs that are semantically equivalent, i.e., where the gold output remains unchanged semantically.
 For the retrieval component, we expect retrieved documents $\mathbf{D}_{i}=\mathcal{R}(x_{i},\mathcal{C})$ and $\mathbf{D}_{i}^{\prime}=\mathcal{R}(x_{i}^{\prime},\mathcal{C})$ to both contain the answer $y_{i}$.
@@
 Formality.
-Formality in language lacks universal definition (Pavlick & Tetreault, 2016; Mosquera & Moreda, 2011; Fang & Cao, 2009), but encompasses situational factors (Hovy, 1987; Lahiri et al., 2011), grammar quality (Peterson et al., 2011), and specific linguistic elements like contractions (Heylighen et al., 1999).
-We quantify formality using the RoBERTa-based formality classifier from (Babakov et al., 2023b).
+Formality in language lacks universal definition (Pavlick and Tetreault, 2016; Mosquera and Moreda, 2011; Fang and Cao, 2009), but encompasses situational factors (Hovy, 1987; Lahiri et al., 2011), grammar quality (Peterson et al., 2011), and specific linguistic elements like contractions (Heylighen and Dewaele, 1999).
+We quantify formality using the RoBERTa-based formality classifier from (Babakov et al., 2023).
 Readability.
-Readability quantifies text comprehensibility through complexity and linguistic features.
+Readability quantifies text comprehensibility through linguistic complexity.
 We employ the widely-used Flesch Reading Ease Score (FRES; Flesch 1948) to assess readability (Rawte et al., 2023; Han et al., 2024), defined in A.1.
@@
 We alter the grammatical correctness through two approaches, inspired by Yin et al.
-(2020); Zhuang & Zuccon (2021); Lichtarge et al.
+(2020); Zhuang and Zuccon (2021); Lichtarge et al.
 (2019): (1) Typos, where random addition, deletion, or substitution operations are applied at a 20% probability per word, requiring an edit distance of at least 1; and (2) Round-trip translation (RTT) via English-Afrikaans-English using EasyNMT with opus-mt model, requiring the output to not be the exact same as the original query.
 3.2 Query Rewriting
-We systematically reformulate queries $x_{i}\rightarrow x_{i}^{\prime}$ across targeted linguistic dimensions while ensuring all rewrites satisfy dimension-specific thresholds and preserve semantic meaning (as specified in Appendix A.1).
-For each distinct dimension—formality, readability, and politeness—we take $5,000$ original queries and use GPT-4o-mini (OpenAI et al., 2024) to generate rewrites using three different prompts (detailed in Appendix G).
-We design distinct query sets across all datasets to be less formal, less readable, or more polite—directions chosen because the well-curated original datasets contained predominantly formal, readable, and neutral language, allowing our modifications to explore more realistic linguistic variations.
-This creates $15,000$ rewritten samples per dimension for each dataset, over which we average results to account for prompt stochasticity.
-For grammatical correctness, we similarly use $5,000$ original queries but apply deterministic transformations through typo introduction and round-trip translation as described in Section 3.1, creating $5,000$ modified samples for each approach.
-We ensure that grammatically modified queries differ from originals using GLEU and edit distance measurements.
-Example rewrites are shown in Table 1.
-For each individual linguistic dimension and dataset, we construct controlled comparative datasets $\mathcal{D}^{\prime}=\{(x_{1}^{\prime},y_{1}),(x_{2}^{\prime},y_{2}),\ldots\}$.
-To validate our approach, we annotate $100$ queries across all linguistic variations, finding that 97.33% of the rewritten queries preserve the semantic meaning of the original query.
-Details are available in Appendix A.2.
 Linguistic Dimension Example Rewrites
@@
 (PopQA) Rewritten: Hey, do you happen to know who wrote Dolores Claiborne? I’m kinda curious!
-Grammar $\downarrow$: Round-Trip Translation Original: Which company is HMS Blankney produced by?
+Grammar: RTT $\downarrow$ Original: Which company is HMS Blankney produced by?
 (EntityQuestions) Rewritten: What company is producing HMS Blankey?
-Grammar $\downarrow$: Typos Original: when did the japanese river otter become extinct
+Grammar: Typos $\downarrow$ Original: when did the japanese river otter become extinct
 (Natural Questions) Rewritten: when did the japanese river otter ecome extinct
 Table 1: Examples of query rewrites across different linguistic dimensions and datasets. Each pair shows the original query and its rewritten form.
+We systematically reformulate queries $x_{i}\rightarrow x_{i}^{\prime}$ across targeted linguistic dimensions while ensuring all rewrites satisfy dimension-specific thresholds and preserve semantic consistency (Appendix A.1).
+Linguistic reformulation.
+For each distinct dimension (formality, readability, and politeness), we sample $5,000$ original queries and use GPT-4o-mini (OpenAI et al., 2024) to generate rewrites using three different prompts (detailed in Appendix I).
+[^2] To verify that our experimental results are robust to rewriting models, we generate $500$ rewritten queries from the PopQA dataset using Llama-3.1-70B-Instruct.
+Shown in Appendix C, results confirm our findings with GPT-4o-mini.
+We design distinct query sets across all datasets to be less formal, less readable, or more polite—directions chosen because the well-curated original datasets contained predominantly formal, readable, and neutral language, allowing our modifications to explore more realistic linguistic variations.
+This creates $15,000$ rewritten samples per dimension for each dataset, over which we average results to account for prompt stochasticity.
+For grammatical correctness, we similarly use $5,000$ original queries but apply deterministic transformations through typo introduction and round-trip translation as described in Section 3.1, creating $5,000$ modified samples for each approach.
+Example rewrites are shown in Table 1.
+Semantic consistency.
+We define a rewritten query $x_{i}^{\prime}$ as semantically consistent with the original query $x_{i}$ if it preserves the original pragmatic intent, i.e., a human annotator would assign the same gold answer $y_{i}$ to both queries.
+This operational definition reflects the requirements that linguistic variation should not alter the underlying information need.
+To strictly enforce this, we apply a two-stage verification procedure.
+First, we automatically filter rewrites using a sentence-level semantic similarity threshold (>0.7>0.7) computed with MPNet-v2 (Song et al., 2020).
+Second, we conduct qualitative evaluation through manual annotation on $250$ queries across all linguistic variations, finding that 94.67% of the rewritten queries preserve the original answer.
+Further details of the annotation are provided in Appendix A.2.
+For each linguistic dimension and dataset, we construct controlled comparative datasets $\mathcal{D}^{\prime}=\{(x_{1}^{\prime},y_{1}),(x_{2}^{\prime},y_{2}),\ldots\}$, enabling direct measurement of performance changes attributable solely to linguistic variation.
 4 Experimental Setups
@@
 MS MARCO (Bajaj et al., 2018) contains questions derived from real user search queries from Bing’s search logs.
-Natural Questions (Kwiatkowski et al., 2019a) contains questions consisting of real, anonymized, aggregated queries to the Google search engine.
+Natural Questions (Kwiatkowski et al., 2019) contains questions consisting of real, anonymized, aggregated queries to the Google search engine.
 Both PopQA and EntityQuestions consist of well-structured, standardized, and simple queries, while queries from MS MARCO and Natural Questions exhibit free-form and arbitrary.
 We evaluate using PopQA’s test split, EntityQuestions’ dev split, and both dev and test splits from MS MARCO and Natural Questions.
+The retrieval is performed on the Wikipedia passage set used in DPR [^3] https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz for the PopQA, EntityQuestions, and Natural Questions datasets, while the MS MARCO dataset uses its corresponding passage dataset (Bajaj et al., 2018).
 4.2 Models
@@
 Generation.
-We evaluate nine most advanced open-source LLMs from three model families with various scales: Llama 3.1 (Llama-3.1-8B-Instruct, Llama-3.1-70B-InstructGrattafiori et al.
-2024), Qwen 2.5 (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct;  Yang et al.
-2024; Team 2024), and Gemma 2 (gemma-2-2b-it, gemma-2-9b-it, gemma-2-27b-itTeam et al.
-2024).
+We evaluate the nine most advanced open-source instruction-tuned LLMs from three model families with various scales: Llama 3.1 (Grattafiori et al.
+2024; 8 and 70 billion), Qwen 2.5Yang et al.
+2024; Team 2024; 3, 7, 32, and 72 billion), and Gemma 2Team et al.
+2024; 2, 9, and 27 billion).
 Detailed hyperparameter settings are included in the Appendix B.1.
@@
 We also include the top five retrieved passages in the context.
-Detailed prompts are provided in Appendix H.
+Detailed prompts are provided in Appendix J.
 4.3 Metrics
@@
 The generation stage is assessed using a comprehensive set of metrics: Answer Match (AM) measures the percentage of the predictions that any substring of the prediction is an exact match of any of the ground truth answers, Exact Match (EM) measures the percentage of predictions that exactly match any of the ground truth answers, and F1 Score (F1) captures the harmonic mean of precision and recall in generated responses.
+We mainly report AM scores in the main paper because it better reflects answer correctness under linguistic and stylistic variaion.
+The full results could be found in Appendix F.
+As a complement to standard metrics, we also use LLM-as-a-Judge with GPT-5-mini (OpenAI, 2025).
 5 RAG Robustness Experimental Results
@@
 We conduct a comprehensive analysis of retrieval systems performance across two candidate retrievers: Contriever and ModernBERT Embed. The results are shown in Table 2.
-Query variations based on linguistic dimensions degrade retrieval performance.
-Our analysis (Table 2) reveals significant linguistic fragility in retrieval systems, with performance degradation averaging 16.7% for Contriever and 13.3% for ModernBERT across all modifications.
-Results show highest sensitivity on PopQA (19.78% average impact), particularly to grammatical transformations (29.34% from RTT).
-MS MARCO exhibits the second-highest impact (16.78%), with striking sensitivity to formality changes (40.41% with Contriever), suggesting retrieval systems may be implicitly optimized for specific linguistic patterns, limiting effectiveness when handling diverse query variations.
 Linguistics Retriever PopQA Entity MARCO NQ $\Delta$ Q-len
@@
 Query linguistic variations degrades retrieval performance consistently across all linguistic characteristics.
-Grammatical variations have the highest impact on retrieval performance.
-On average, grammaticality rewrites emerge as the most impactful linguistic variation on recall performance.
-Round-trip translation degrades recall by an average of 19.56% across all datasets.
-Interestingly, ModernBERT shows greater vulnerability to these structural transformations (20.12% drop) compared to Contriever (19% drop).
-Typographical errors present another significant challenge, causing an average recall reduction of 18.51%.
-However, the retrievers display opposite behavior patterns with typos: Contriever exhibits substantially lower robustness (22.22% drop) than ModernBERT (14.81% drop).
-This suggests that ModernBERT’s diverse training data mixture likely enables it to develop greater robustness to character-level grammatical perturbations compared to Contriever.
-Politeness variations have minimal impact on retrieval performance.
-Politeness variations have the least impact on retrieval performance, with an average recall drop of only 6.48% across all datasets and retrievers.
-This stands in stark contrast to grammatical variations (19.56%) and typos (18.51%).
-The minimal effect is most evident in Natural Questions with Contriever (1.16%) and EntityQuestions with Contriever (1.70%).
-This suggests that retrieval models effectively filter out social courtesy markers while preserving their focus on the query’s core semantic content and keywords, maintaining robust performance despite changes in query politeness level.
-Retrieval performance drops independent of query length.
-Our analysis, as shown in Table 2 demonstrates that query length changes do not directly correlate with retrieval performance.
-Queries with increased formality showed substantial length increases (+13.65 tokens) yet produced inconsistent performance impacts across datasets.
-Conversely, round-trip translated queries were marginally shorter (-0.26 tokens) but consistently caused significant performance degradation.
-This indicates retrieval models respond more to linguistic quality (grammatical correctness, readability) than to query length itself, highlighting the need for systems robust to linguistic variations rather than optimized for specific query lengths.
-Scaling up number of documents improves performance.
-Table 2 shows performance degradation ($\Delta$R@K) decreasing as $K$ increases, indicating linguistic perturbations cause relevant documents to slide down rather than disappear from the ranked list.
-As an example, for rewrites with typos, $\Delta$ R@K for Contriever decreases from 22.2 at R@5 to 12.0 at R@100, and for ModernBERT from 20.1 to 9.8.
-This is detailed further in Appendix C.1.
-This ranking deterioration forces downstream language models to operate with suboptimal information, potentially compromising response quality.
-We further investigate this hypothesis in Section 5.4.2 by examining if rerankers can improve recall scores for Top-5 retrieved documents.
-5.2 Generation Analysis
-The RAG experiment results on ModernBERT retriever and nine LLMs are presented in Table 3. Overall, RAG systems show performance degradation on all linguistic variations.
+Query variations based on linguistic dimensions degrade retrieval performance.
+Our analysis (Table 2) reveals significant linguistic fragility in retrieval systems, with performance degradation averaging 16.7% for Contriever and 13.3% for ModernBERT across all modifications.
+We hypothesize that ModernBERT’s increased robustness could largely be attributed to its diverse pre-training data, which forces the model to learn to process non-standard sequences and enhances its resilience to linguistic noise.
+Additionally, the masked language modeling objective utilized by ModernBERT could effectively train the model to denoise and reconstruct context from corrupted inputs.
+Results show highest sensitivity on PopQA (19.78% average impact), particularly to grammatical transformations (29.34% from RTT).
+MS MARCO exhibits the second-highest impact (16.78%), with striking sensitivity to formality changes (40.41% with Contriever), suggesting retrieval systems may be implicitly optimized for specific linguistic patterns, limiting effectiveness when handling diverse query variations.
 Model Rew. $\downarrow$ (Ori.) Readability Grammatical Correctness
@@
 All systems exhibit performance drops across all linguistic variations and datasets.
+Grammatical variations have the highest impact on retrieval performance.
+On average, grammaticality rewrites emerge as the most impactful linguistic variation on recall performance.
+Round-trip translation degrades recall by an average of 19.56% across all datasets.
+Interestingly, ModernBERT shows greater vulnerability to these structural transformations (20.12% drop) compared to Contriever (19% drop).
+Typographical errors present another significant challenge, causing an average recall reduction of 18.51%.
+However, the retrievers display opposite behavior patterns with typos: Contriever exhibits substantially lower robustness (22.22% drop) than ModernBERT (14.81% drop).
+This suggests that ModernBERT’s diverse training data mixture likely enables it to develop greater robustness to character-level grammatical perturbations compared to Contriever.
+Politeness variations have minimal impact on retrieval performance.
+Politeness variations have the least impact on retrieval performance, with an average recall drop of only 6.48% across all datasets and retrievers.
+This stands in stark contrast to grammatical variations (19.56%) and typos (18.51%).
+The minimal effect is most evident in Natural Questions with Contriever (1.16%) and EntityQuestions with Contriever (1.70%).
+This suggests that retrieval models effectively filter out social courtesy markers while preserving their focus on the query’s core semantic content and keywords, maintaining robust performance despite changes in query politeness level.
+Retrieval performance drops independent of query length.
+Our analysis, as shown in Table 2 demonstrates that query length changes do not directly correlate with retrieval performance.
+Queries with increased formality showed substantial length increases (+13.65 tokens) yet produced inconsistent performance impacts across datasets.
+Conversely, round-trip translated queries were marginally shorter (-0.26 tokens) but consistently caused significant performance degradation.
+This indicates retrieval models respond more to linguistic quality (grammatical correctness, readability) than to query length itself, highlighting the need for systems robust to linguistic variations rather than optimized for specific query lengths.
+We confirm this by analyzing semantic preservation using LLM-as-a-judge in Table 4 and a human evaluation of results in Table 5.
+Scaling up number of documents improves performance.
+[…]
+Dataset Readability Formality Politeness Gram-RTT Gram-Typos
+PopQA 19.41% 11.79% 6.75% 35.60% 21.20%
+NQ 13.12% 4.27% 4.38% 22.74% 5.63%
+MARCO 10.56% 4.05% 1.75% 15.21% 5.32%
+Entity 13.20% 5.06% 2.86% 22.18% 10.49%
+Table 4: RAG performance in LLM-as-a-Judge evaluation (ModernBERT + Qwen2.5-72B-Instruct), with relative percentage performance degradation.
+LLM-as-A-Judge evaluation shows a consistent degradation pattern.
+5.2 Generation Analysis
+The RAG experiment results on ModernBERT retriever and nine LLMs are presented in Table 3 with answer match (AM) scores.
+Table 4 shows the results with LLM-as-a-Judge evaluation.
+Overall, RAG systems show performance degradation on all linguistic variations.
 RAG systems are sensitive to linguistic variations.
@@
 These findings suggest that while the RAG systems perform well on standard NLP benchmarks with structured queries, they remain vulnerable to common linguistic variations.
-The full experiment results are shown in Appendix D.
+The full experiment results are shown in Appendix F.
 Politeness reformulations yield different impacts on AM and EM scores.
@@
 Specifically, we observe 44.57% and 18.32% drops in EM scores for queries from the Natural Questions and PopQA datasets, respectively.
-Through manual checks, we find that LLMs tend to generate more complete or formal responses in polite query formulations, which results in lower EM rates but similar performance on AM.
+Although the rewritten queries preserve the original pragmatic intent, they introduce longer phrasing and auxiliary words that shift embeddings and sometimes nudge generators toward more verbose outputs.
+As a result, they still cause slight regressions in the RAG.
 The round-trip translation errors and typos highlight different sensitivities.
@@
 This counterintuitive finding may be attributed to the structural transformations introduced during translation that become more problematic for larger models attempting more precise reasoning.
+Human annotator evaluation aligns with our automated evaluation metrics
+To validate our automated metrics and control for potential confounds such as length or verbosity, two independent annotators evaluated 250 samples across EntityQuestions (100 samples) and NaturalQuestions (150 samples).
+For typos and RTT variations, rewrites with more than 2 character edits to any entity were marked as not preserving semantic meaning.
+As shown in Table 5, semantic preservation rates ranged from 86.7% to 100% with high inter-annotator agreement (96.7%-100%), confirming that observed performance differences stem from linguistic variation rather than semantic drift.
+Full annotation protocols are provided in Section A.2.
 5.4 Exploring the Robustness of Advanced RAG Systems
+Many modern RAG systems include more components than simply retrieval and generation, which aim to make them more useful for users (Gao et al., 2022).
+In this section, we explore the possibility of using a simple query-expansion step to fix the vulnerability on linguistic variations and whether the addition of reranking improves RAG robustness.
+Detailed results of the experiment can be found in Appendices G and H.
 Figure 3: Retrieval (ModernBERT, R@5 Score) and generation (Qwen2.5-7B-Instruct, AM Score) performance across different RAG settings on PopQA.
@@
 (3) Reranking improves performance on both original and rewritten queries.
-Many modern RAG systems include more components than simply retrieval and generation which aim to make them more useful for users (Gao et al., 2022).
-In our experiments, we broadly explore whether the addition of two advanced components—Query Expansion and Reranking—improves RAG robustness.
-Detailed results of the experiment can be found in Appendices E and F.
 5.4.1 Query Expansion
@@
 To address this, we incorporate a Cross-Encoder-based re-ranker to significantly enhance the quality of final answers.
-Specifically, we employ the MS MARCO Cross-Encoders developed by Reimers & Gurevych (2019) to re-rank passages retrieved by ModernBERT and Contriever.
+Specifically, we employ the MS MARCO Cross-Encoders developed by Reimers and Gurevych (2019) to re-rank passages retrieved by ModernBERT and Contriever.
 As illustrated in Figure 3, re-ranking substantially improves retrieval performance, particularly for rewritten queries, achieving an average improvement of 16.56% compared to only 7.40% for original queries.
@@
 These findings highlight the need to develop strategies that ensure reliable performance across linguistically varied queries, guiding future improvements of real-world RAG systems.
+Limitations
+While our choice of linguistic dimensions cover a broad spectrum of stylistic, pragmatic, and structural variations, other relevant factors such as dialect, idiomatic expressions, or domain-specific terminology could be explored in future work.
+Our grammatical variations (typos and round-trip translation) represent a subset of potential errors; other categories such as verb form or agreement errors Dahlmeier et al.
+(2013) may present greater challenges and warrant future investigation.
+We conducted query rewriting using two LLMs (GPT-4o-mini and Llama-3.1-70B-Instruct) and observed similar vulnerabilities; future studies may verify the generalizability of these findings using a broader range of rewriting methods.
+Additionally, while we explored widely used methods such as query expansion and reranking to test for mitigation strategies, more comprehensive approaches, including training models explicitly on diverse, linguistically varied data, remain important avenues for future research.
+Acknowledgments
+This research was supported in part by the National Science Foundation under grant 2230466 and in part by DSO National Laboratories.
 Appendix A Linguistic Rewriting Settings
@@
 Therefore, in the query-rewriting process, we add an additional constraint where the sentence similarity between the original and rewritten queries must be greater than 0.7, as assessed by the MPNet-v2 model Song et al.
-(2020) in the SentenceBert Library Reimers & Gurevych (2019).
+(2020) in the SentenceBert Library Reimers and Gurevych (2019).
 Evaluation
 : Flesch Reading score is defined as follows:
-FRES=206.8351.015(total wordstotal sentences)84.6(total syllablestotal $\text{FRES}=206.835-1.015\left(\frac{\text{total words}}{\text{total sentences}}\right)-84.6\left(\frac{\text{total syllables}}{\text{total words}}\right)$ (1)
+$\begin{split}\mathrm{FRES}&=206.835-1.015\;\frac{\text{total words}}{\text{total sentences}}\\ &\quad-84.6\;\frac{\text{total syllables}}{\text{total words}}\end{split}$ (1)
 A.2 Semantic Preservation of Linguistic Variations
-To ensure our linguistic rewrites preserve semantic meaning, we conducted a qualitative evaluation where two researchers annotated 20 randomly sampled query rewrites per linguistic property for the EntityQuestions dataset, totaling 100 queries examined by both annotators.
-Our primary assessment criterion was whether the rewritten question remained semantically similar enough to warrant the same answer as the original query.
-For typos and back-translation variations specifically, we established a threshold of maximum 2 character edits on entities in the questions to determine if a query remained semantically equivalent to its original form.
+To ensure our linguistic rewrites preserve semantic meaning, we conducted a qualitative evaluation where two independent researchers annotated randomly sampled query rewrites per linguistic property from EntityQuestions (100 examples) and Natural Questions (150 examples) datasets, totaling 250 queries.
+A.2.1 Annotation Guidelines
+Annotators were provided with the following protocol to assess whether each rewritten query preserved the semantic meaning of the original:
+  • Mark as semantically preserved if both queries would elicit identical factual answers
+  • Mark as not preserved if the rewrite changes entities, relationships, or temporal/spatial constraints
+  • Rule for Typos and RTT: If more than 2 characters are edited in any entity mention, mark as not preserved.
 This threshold helps maintain a balance between introducing realistic linguistic variations and preserving essential meaning.
-Our analysis revealed high semantic preservation rates across all linguistic dimensions:
-Linguistic Dimension Semantic Preservation (%) Inter-annotator Agreement (%)
+  • Ignore stylistic differences that do not affect the core information need
+Annotators worked independently and their judgments were compared to compute inter-annotator agreement.
+Linguistic Variation Semantic Inter-annotator
+Preservation (%) Agreement (%)
 Politeness 100.0 100.0
@@
 Typos 95.0 100.0
-Round-trip Translation 95.0 100.0
-Table 4: Percentage of queries that maintain semantic similarity after linguistic rewrites and inter-annotator agreement rates.
+RTT 95.0 100.0
+(a) EntityQuestions
+Linguistic Variation Semantic Inter-annotator
+Preservation (%) Agreement (%)
+Politeness 100.0 100.0
+Formality 90.0 96.7
+Readability 100.0 100.0
+Typos 86.7 96.7
+RTT 96.7 96.7
+(b) Natural Questions
+Table 5: Percentage of queries that preserve semantic meaning after linguistic rewrites, along with inter-annotator agreement rates across 250 annotated samples.
+A.2.2 Results and Analysis
+Table 5 shows high semantic preservation rates across all linguistic dimensions, ranging from 86.7% to 100%.
+The high inter-annotator agreement (96.7%-100%) further validates the reliability of our assessment approach.
 These results demonstrate the robustness of our rewriting methodology in preserving semantic content while introducing natural linguistic variations.
-The high inter-annotator agreement further validates the reliability of our assessment approach.
+Notably, even with our strict 2-character edit threshold for typos and RTT, we achieve high preservation rates (86.7%-96.7%), indicating that our generation process successfully balances realism with semantic fidelity.
 This semantic preservation is critical for our experimental design, as it ensures that performance differences observed in RAG systems can be attributed to linguistic variations rather than semantic drift.
@@
 We set temperature as 0.5 and top_q 0.90 to strike a balance between creativity and accuracy. The max_tokens is set to 128 considering the gold answers length.
-Appendix C Full Retrieval Experiment Results
+Appendix C Experiment Results on Llama 3.1 Rewriting Queries
+We conducted supplementary experiments by sampling 500 queries from the PopQA dataset.
+We generated rewritten queries using Llama-3.1-70B-Instruct, employing the same rewriting criteria as previously described.
+Then, we performed RAG experiments using ModernBERT as the retriever and Qwen2.5-7B-Instruct as the generator.
+The results are summarized below:
+Dimension Original Rewritten $\boldsymbol{\Delta}$
+Readability 0.624 0.490 21.5%
+Formality 0.826 0.746 9.7%
+Politeness 0.728 0.626 14.0%
+Table 6: Retrieval Performance (R@5)
+Dimension Original Rewritten $\boldsymbol{\Delta}$
+Readability 0.500 0.372 25.6%
+Formality 0.776 0.684 11.9%
+Politeness 0.604 0.520 13.9%
+Table 7: RAG Generation Performance (Answer Match - AM Score)
+In general, the RAG system remains notably sensitive to the linguistic variations introduced by LLaMA.
+Specifically, queries rewritten for reduced readability caused the most significant performance degradation, with 21.5% in retrieval accuracy and 25.6% in the AM score.
+Unlike variations introduced by GPT, the system exhibited greater sensitivity to changes in politeness compared to formality.
+We will present a comprehensive analysis and more detailed results in the revised paper.
+Appendix D Data and Code Availability
+Our code and rewritten query datasets will be released after the peer review stage under the CC BY 4.0 license.
+All existing datasets, models, and codes used in this work were employed consistently with their intended research purposes.
+Appendix E Full Retrieval Experiment Results
 In this section we provide the absolute results from Contriever retriever experiments that led to Table 2
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 EntityQuestions RTT 0.6838 0.7332 0.7710 0.8422
@@
 Table 8: Contriever Retrieval performance (R@k) of original queries across datasets and linguistic modifications.
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 EntityQuestions RTT 0.5842 0.6428 0.6926 0.7896
@@
 Table 9: Contriever Retrieval performance (R@k) of rewritten queries across datasets and linguistic modifications
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 EntityQuestions RTT 0.6614 0.7184 0.7598 0.8284
@@
 Table 11: ModernBERT Retrieval performance (R@k) for rewritten queries across datasets and linguistic modifications
-C.1 Scaling Number of documents
+E.1 Scaling Number of documents
 Figure 4: Average Recall@K increase as Number of Top-K Documents increases – Contriever.
@@
 This hypothesis is confirmed in Section 5.4.2, where the reranking shows considerable performance gains at R@5, showing that retrieval systems tend to push the correct documents for rewritten queries to lower ranks, and reranking helps prioritize them again, which is demonstrated by the reduction in performance degradation in Figure 3.
-Appendix D Full RAG Experiment Results
+Appendix F Full RAG Experiment Results
 Contriever, Formality, AM Score
@@
 Table 17: RAG experiment results with ModernBERT as retrieval model on F1 scores.
-Appendix E Full Query Expansion Experiment Results
+Appendix G Full Query Expansion Experiment Results
 The following tables provide results of Retrieval and Generation components across MS MARCO and PopQA.
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 MS MARCO RTT 37.10 43.74 50.56 63.76
@@
 Table 18: Contriever retrieval results with RAG+HyDE for original queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 MS MARCO RTT 32.80 38.74 44.46 57.70
@@
 Table 19: Contriever retrieval results with RAG+HyDE for rewritten queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 PopQA RTT 0.5984 0.6496 0.6980 0.7936
@@
 Table 20: ModernBERT retrieval results with RAG+HyDE for original queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 PopQA RTT 0.3750 0.4230 0.4678 0.5802
@@
 Table 27: RAG experiment results with query expansion and ModernBERT as retrieval model on F1 scores.
-Appendix F Full Reranking Experiment Results
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Appendix H Full Reranking Experiment Results
+Dataset Linguistics R@5 R@10 R@20 R@100
 MS MARCO RTT 0.4620 0.5208 0.5618 0.6102
@@
 Table 28: Contriever retrieval results with RAG+Reranking for original queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 MS MARCO RTT 0.4144 0.4674 0.5158 0.5754
@@
 Table 29: Contriever retrieval results with RAG+Reranking for rewritten queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 PopQA RTT 0.6836 0.7462 0.7934 0.8344
@@
 Table 30: ModernBERT retrieval results with RAG+Reranking for original queries
-Dataset Linguistic Dimension R@5 R@10 R@20 R@100
+Dataset Linguistics R@5 R@10 R@20 R@100
 PopQA RTT 0.5096 0.5744 0.6204 0.6856
@@
 Table 37: Generation results with ModernBERT retrieval and reranking on F1 scores.
-Appendix G Rewriting Prompts
-G.1 Formality
+Appendix I Rewriting Prompts
+I.1 Formality
 Prompt 1:
@@
 Original Query:
-G.2 Readablility
+I.2 Readablility
 Prompt 1:
@@
 Less Readable Query:
-G.3 Politeness
+I.3 Politeness
 Prompt 1:
@@
 Polite Query:
-Appendix H LLMs Prompts
-H.1 Few-shot Prompts
+Appendix J LLMs Prompts
+J.1 Few-shot Prompts
 Few-shot examples:
@@
 Now, it’s your turn to answer the question below. The answer should contain ONLY one sentence and DO NOT explain reasons.
-Appendix I Rewriting Examples
+Appendix K Rewriting Examples
 Category Type Query
@@
 Table 38: Rewriting examples across all linguistic variations from the EntityQuestions dataset, with queries split across rows for readability.
+Appendix L Computational Resources
+Our experimental setup utilized models of varying scales: Gemma-2 (2B, 9B, 27B parameters), Llama-3.1 (8B, 70B parameters), and Qwen-2.5 (3B, 7B, 32B, 72B parameters).
+For retrieval, we employed ModernBERT Embed (149M parameters) and Contriever.
+We conducted comprehensive evaluations across 5 linguistic dimensions (4 dimensions plus 2 grammatical correctness subtypes), 4 datasets, 9 language models, and 2 retrieval systems, totaling 360 experimental configurations.
+Each model inference run required approximately 1.5 hours, resulting in 540 GPU hours on 16 L40S GPUs distributed across different model configurations.
+Retrieval evaluation required an additional 40 GPU hours.
+Data rewriting was performed using GPT-4o-mini, requiring 40 hours of API usage.
+Total computational cost comprised 620 GPU hours on L40S hardware plus commercial API usage for data preprocessing.